### v1

# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

Abstract
Accurate and efficient skin lesion classification on edge devices is critical for accessible dermatological care but remains challenging due to computational, energy, and privacy constraints. We introduce QANA, a novel quantization-aware neuromorphic architecture for incremental skin lesion classification on resource-limited hardware. QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference. Its quantization-aware head and spike-compatible transformations enable seamless conversion to spiking neural networks (SNNs) and deployment on neuromorphic platforms. Evaluation on the large-scale HAM10000 benchmark and a real-world clinical dataset shows that QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% Top-1 accuracy and 81.7% macro F1 on the clinical dataset, consistently outperforming leading CNN-to-SNN models under fair comparison. Deployed on BrainChip Akida hardware, QANA achieves 1.5 ms inference latency and 1.7 mJ energy per image, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, and exceeding the performance of advanced CNN-to-SNN conversion methods. These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
Index Terms:
Resource-Constrained Devices, Edge Computing, Neuromorphic Computing, Low Latency, Energy Efficiency.
I Introduction
Skin diseases present significant diagnostic challenges, particularly for clinicians without specialized training, leading to frequent misdiagnosis of conditions such as melanoma and Merkel cell carcinoma [1, 2]. Although deep learning-based diagnostic systems have shown promising performance [3], most approaches rely on centralized training and inference, requiring sensitive patient data to be transferred to cloud servers [4]. This raises data security risks [5] and is constrained by strict privacy regulations such as HIPAA [6] and GDPR [7]. Furthermore, to support dermatological care in home and remote settings lacking conventional healthcare infrastructure[8], it is essential to develop models that enable on-device training and inference.
However, deploying deep learning models on edge devices faces several challenges. Conventional CNNs require significant computational and energy resources, increasing device complexity and limiting portability [9]. In addition, effective training typically depends on large labeled datasets, which are often unavailable for rare skin diseases [10, 11]. Frequent model updates further introduce computational overhead and privacy concerns, while hardware constraints restrict model complexity and inference accuracy on edge platforms [2].
Figure 1: Detailed architecture of our end-to-end framework for quantization-aware neuromorphic skin lesion classification: (1) data preprocessing (quality filtering, augmentation, and SMOTE-based oversampling); (2) a novel quantization-aware network for feature extraction and spike-compatible transformation; (3) CNN-to-SNN conversion with operator mapping and temporal spike encoding; and (4) SNN deployment with on-chip optimization for real-time and energy-efficient inference on edge hardware.
Recently, Spiking Neural Networks (SNNs) and neuromorphic computing platforms have emerged as promising alternatives for overcoming deployment challenges in dermatological diagnosis [12]. Unlike conventional CNNs, SNNs use discrete spike events for information encoding and transmission, resulting in sparse and event-driven computation that greatly reduces power consumption [13]. This spike-based encoding is well-suited for scenarios with limited labeled data, as the temporal dynamics of spikes facilitate effective learning from fewer examples [14]. Neuromorphic processors, such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17], natively support SNNs and enable on-chip incremental learning. Incremental learning enables these systems to adapt efficiently to new patient data without full model retraining, aligning well with clinical practices that regularly acquire new diagnostic cases [18]. Additionally, these neuromorphic devices possess compact physical footprints and substantially lower energy requirements compared to conventional GPUs [19], making them viable candidates for edge-based deployment in portable diagnostic instruments.
The prevalent approach for neuromorphic deployment is to first train conventional CNN architectures, such as ResNet or DenseNet, on large datasets, and then convert them to equivalent SNN models for edge inference [20]. While this conversion aims to combine the accuracy of CNNs with the efficiency of SNNs [21], several practical limitations persist. Key CNN components, such as Batch Norm, Flatten, and Global Average Pooling, cannot be directly mapped to spike-based neuron units, complicating and often degrading the conversion process [22]. In addition, converted SNNs frequently experience significant accuracy loss on small, imbalanced medical datasets due to quantization effects and limited capacity for capturing subtle pathological features [23].
In this work, we present an end-to-end pipeline for efficient and incremental skin lesion classification on neuromorphic hardware. As illustrated in Fig. 1, our approach consists of four main stages: (1) data preprocessing, including image quality screening, augmentation, and SMOTE-based class balancing; (2) construction of a quantization-aware neural network with stacked Ghost modules, channel attention, and a spike-compatible output head; (3) automated conversion of the trained network to a spiking neural network (SNN) using the Akida MetaTF toolkit; and (4) deployment and optimization on the BrainChip Akida neuromorphic platform for low-latency, energy-efficient inference. The core QANA architecture is specifically designed for hardware compatibility, enabling direct and lossless translation to spike-based event-driven computation. This pipeline directly addresses the major challenges of edge deployment, including computational constraints, class imbalance, limited data, and the need for incremental model updates.
Experimental results on both the HAM10000 public benchmark and a real-world clinical dataset demonstrate the practical advantages of our approach. QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% / 81.7% on the clinical dataset. On Akida hardware, the system delivers 1.5 ms inference latency and 1.7 mJ energy consumption per image, surpassing state-of-the-art CNN-to-SNN conversion baselines in both accuracy and efficiency. These findings confirm the effectiveness of QANA for resource-constrained, real-time medical image analysis.
II Related Work
Prior studies on neuromorphic hardware and Spiking Neural Networks (SNNs) have highlighted their suitability for edge inference and continual learning in constrained environments [14]. Platforms such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17] demonstrate ultra-low-power operation and support on-chip incremental learning, enabling model updates without full retraining. These features are advantageous for medical imaging tasks on edge devices. However, existing works primarily focus on general benchmarks like MNIST [24] or CIFAR [25] and rarely address domain-specific challenges in dermatology [26]. Furthermore, in medical image contexts, edge Spiking Neural learning methods often suffer from catastrophic forgetting and fail to preserve performance on previously learned classes when new samples are introduced [27].
CNN-to-SNN conversion has become a common strategy to leverage mature CNN architectures—ResNet, DenseNet—for deployment on neuromorphic processors [28]. Conversion toolkits, including Akida’s CNN2SNN [29], deliver compatibility by quantizing weights and replacing activation functions with spike-based equivalents. While some studies report preserved accuracy on standard datasets [30], performance significantly degrades on small or imbalanced medical datasets due to batching overfitting, quantization noise, and inability to capture rare lesion features [15]. Limitations in mapping CNN operations—such as batch normalization, global average pooling, and multi-bit activations—to spiking neurons further compound accuracy loss during conversion [31].
Figure 2: Detailed architecture of our QANA, which performs iterative feature extraction using stacked Ghost modules, ECA, and residual blocks, followed by spike-compatible transformation with batch normalization, ReLU activation, and Squeeze-and-Excitation (SE) block. The output is then quantized and projected to class logits for SNN deployment.
In parallel, conventional CNN-based methods for skin lesion classification, such as those using DenseNet-121 [32] or Inception-v4 [33], show high performance on benchmark datasets like ISIC. Despite their effectiveness, these models necessitate continuous retraining on cloud infrastructure to address domain shifts, new lesion types, or demographic variations [18]. Such retraining incurs latency, high computational cost, and data privacy concerns [34]. Moreover, general CNN models often fail to generalize to underrepresented conditions [35], such as rare tumors or images from diverse skin tones, highlighting the need for adaptive and privacy-preserving on-device solutions. These limitations motivate the development of neuromorphic frameworks that maintain diagnostic performance while supporting incremental learning and efficient operation in resource-constrained settings.
III Methodology
This section details the complete pipeline for the development and deployment of the proposed neuromorphic skin lesion classification system. As shown in Fig. 1, the pipeline encompasses data preprocessing, the design of a quantization-aware neural network architecture compatible with spiking inference, conversion to an event-driven SNN format, and hardware-level deployment on the Akida platform.
III-A Data Preprocessing
The pipeline includes three main stages: image preprocessing to ensure quality and compatibility with the Akida hardware, data augmentation to enhance dataset diversity and robustness, and SMOTE to address severe class imbalance among lesion categories.
III-A1 Image Preprocessing
All dermatoscopic images are first screened for quality. Corrupted, low-resolution, or artifact-contaminated images are excluded using automated checks based on image metadata and pixel statistics, followed by manual review when ambiguity remains. For compatibility with the Akida neuromorphic platform, all accepted images are resized to $64\times 64$ pixels using bilinear interpolation, which is the maximum supported resolution under current on-chip memory constraints. Pixel intensities are normalized channel-wise to the range $[0,1]$ to ensure consistent input distribution. No denoising or inpainting operations are performed, as quality control steps eliminate samples with significant artifacts.
III-A2 Data Augmentation
To mitigate overfitting and increase the diversity of the training dataset, each image in the training split undergoes stochastic augmentation with the following operations: brightness adjustment (random factor sampled uniformly from $[0.7,1.3]$), contrast modification ($[0.8,1.2]$), random horizontal and vertical flipping (each with probability 0.5), hue shift (random shift in $[-0.08,0.08]$ in HSV space), and saturation variation ($[0.85,1.15]$). Each augmentation is applied independently to each sample with the specified probability. All augmentation parameters are determined by fixed random seeds for reproducibility. These transformations simulate real-world imaging variation while preserving lesion semantics.
III-A3 Synthetic Minority Oversampling (SMOTE)
To address class imbalance, especially in rare lesion categories, the Synthetic Minority Oversampling Technique (SMOTE) [36] is applied to the training data. Given a minority class dataset $\mathcal{S}=\{x_{1},x_{2},...,x_{n}\}$ in $\mathbb{R}^{d}$, for each $x_{i}\in\mathcal{S}$, the $k$ nearest neighbors $\mathcal{N}_{i}=\{x_{i_{1}},...,x_{i_{k}}\}$ are identified in feature space via Euclidean distance. For each synthetic sample, a random neighbor $x_{i_{j}}$ is selected, and a new sample is generated by:
$x_{\text{new}}=x_{i}+\lambda\cdot(x_{i_{j}}-x_{i}),\quad\lambda\sim\mathcal{U}(0,1)$ (1)
where $\lambda$ is independently sampled for each feature dimension. Repeating this process for all $x_{i}$ and multiple neighbors results in the augmented set:
$\mathcal{S}_{\text{aug}}=\mathcal{S}\cup\left\{x_{i}+\lambda^{(m)}\cdot(x_{i_{j_{m}}}-x_{i})\mid x_{i}\in\mathcal{S},\,1\leq m\leq M\right\}$ (2)
where $M$ is the number of synthetic samples generated per original sample. All images are maintained at $64\times 64$ resolution to conform with Akida’s memory requirements throughout oversampling. SMOTE is applied only to the training set to preserve test set integrity and eliminate information leakage.
III-B Quantization-Aware Network Architecture
The model (as shown in Fig. 2) comprises a hierarchical cascade of Ghost-based multi-scale feature extraction blocks, integrated channel attention mechanisms, and a quantization-aware transformation stage that produces SNN-ready outputs. This architecture balances computational efficiency and discriminative power, facilitating direct conversion and robust inference on the Akida neuromorphic platform.
III-B1 Iterative Feature Extraction and Downsampling
Feature extraction begins with a stack of Ghost blocks [37], designed to maximize representational diversity while minimizing computational cost. As shown in Fig. 3, each Ghost block receives a feature tensor $F^{(l-1)}$ (for $l=1,\ldots,4$) and applies both base and ghost convolutions as:
$F_{\text{ghost}}^{(l)}=\text{Concat}\Big{(}$ $\underbrace{\mathcal{F}_{1\times 1}^{(l)}\big{(}F^{(l-1)}\big{)}}_{\text{base: pointwise conv}},\,\underbrace{\mathcal{F}_{3\times 3}^{(l)}\big{(}F^{(l-1)}\big{)}\odot\mathbf{M}^{(l)}}_{\text{ghost: depthwise conv with filter mask}}\Big{)}$ (3)
Here, $\mathcal{F}_{k\times k}^{(l)}$ is a separable convolution of kernel size $k\times k$, and $\mathbf{M}^{(l)}$ is a binary mask selecting learnable ghost filters. This combination generates both local and extended receptive fields, enabling richer features with fewer FLOPs.
After concatenation, the tensor is processed by batch normalization, quantization-bounded activation, and dropout:
$F_{\text{drop}}^{(l)}=\text{Dropout}\Big{(}\min\big{(}6,\,\max(0,\,\gamma^{(l)}\cdot\text{BN}(F_{\text{ghost}}^{(l)})+\beta^{(l)})\big{)}\Big{)}$ (4)
This quantization-aware step is crucial for downstream SNN compatibility and robust generalization.
Figure 3: Schematic of the Ghost module. The input feature map is first processed by a lightweight convolution to extract a reduced set of primary features with channel size $\mu C$, where $C$ is the target output dimensionality and $\mu\in(0,1)$ is a tunable ratio. Subsequently, inexpensive operations are applied to the primary features to generate additional ghost features of size $(1-\mu)C$. These are concatenated along the channel axis to form the final output of size $C$.
To capture cross-channel correlations, each block integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) mechanism, inspired by Efficient Channel Attention (ECA) mechanism [38]. As shown in Fig. 4, Instead of computing global channel statistics via global average pooling and 1D convolution, we adopt a lightweight depthwise convolution followed by pointwise channel-wise scaling. This enables efficient modeling of spatial-channel dependencies with minimal overhead and neuromorphic compatibility:
$\widetilde{F}^{(l)}=\sigma\left(\mathbf{W}_{1\times 1}^{(l)}*\text{BN}\left(\text{DWConv}_{k\times k}\left(F^{(l)}\right)\right)\right)\odot F^{(l)}$ (5)
where $\text{DWConv}_{k\times k}$ and $\mathbf{W}_{1\times 1}^{(l)}$ jointly form the attention mechanism in our SA-ECA block, and $\sigma$ is the sigmoid activation used to generate the attention mask. This lightweight attention preserves channel expressiveness with minimal overhead.
To ensure stable deep stacking, we employ residual skip connections and spatial downsampling:
$F^{(l)}=\text{MaxPool2D}\left(\boldsymbol{\alpha}^{(l)}\odot F_{\text{drop}}^{(l)}+\mathbf{P}^{(l)}F^{(l-1)}\right)$ (6)
where $\mathbf{P}^{(l)}$ projects the previous block’s output for dimension alignment if needed. This design preserves gradient flow, reduces vanishing/exploding risk, and aggregates multi-scale context, all critical for reliable feature extraction in small-data regimes. After four such blocks, the spatial size is reduced to $4\times 4$.
Figure 4: Illustration of the Spatially-Aware ECA (SA-ECA) block. A depthwise convolution is first applied to extract channel-wise statistics, followed by a lightweight 1D convolution to model local channel dependencies. The resulting attention weights are used to rescale the input feature channels, enhancing discriminative information with minimal computational overhead.
Figure 5: Illustration of the Squeeze-and-Excitation (SE) block. The input feature map undergoes global pooling, followed by two fully connected layers with ReLU and sigmoid activations to compute channel-wise weights. The original feature map is then rescaled by these weights, enabling adaptive recalibration of channel responses.
III-B2 Spike-Compatible Feature Transformation
The output of the previous stage, $F^{(4)}$, is passed to a spike-compatible transformation module engineered for direct quantization and SNN integration. A SeparableConv2D $(3\times 3,256)$ [39] generates higher-dimensional features, followed by quantization-aware normalization and bounded activation:
$\widehat{F}=\min\left(1,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (7)
This mapping guarantees that all activations lie in $[0,1]$, which is both compatible with spike encoding and preserves information for subsequent inference.
To further optimize the channel-wise information flow, as shown in Fig. 5, a Squeeze-and-Excitation (SE) block [40] is applied, implementing a two-stage bottleneck and gating mechanism:
$\mathbf{s}=\sigma\left(W_{2}\,\delta\left(W_{1}\frac{1}{16}\sum_{i,j}\widehat{F}(i,j,:)\right)\right)$ (8)
where $W_{1}$ and $W_{2}$ are dense layers, $\delta$ is ReLU, and $\sigma$ is sigmoid. Each feature channel is then scaled by $s_{c}$, which adaptively modulates discriminative capacity and provides additional regularization for small, imbalanced datasets.
III-B3 Quantized Output Projection
The final block flattens the spike-compatible features $\widehat{F}$ to a vector $r\in\mathbb{R}^{4096}$ and applies a linear projection to produce the model output:
$\mathbf{y}=\mathbf{W}_{\text{cls}}\,r+\mathbf{b}_{\text{cls}}$ (9)
Here, $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{7\times 4096}$, $\mathbf{b}_{\text{cls}}\in\mathbb{R}^{7}$. The output $\mathbf{y}$ is a $7$-dimensional, already quantized vector that can be seamlessly passed to the SNN converter for neuromorphic inference.
III-C CNN-to-SNN Conversion for Neuromorphic Deployment
The conversion of a trained convolutional neural network (CNN) into a spiking neural network (SNN) is a critical step for enabling event-driven inference on neuromorphic hardware. In this work, we utilize the Akida MetaTF toolkit [41, 15] to perform an automated and quantization-aware transformation of the CNN backbone described above into an SNN model suitable for direct deployment on the BrainChip Akida processor.
III-C1 Conversion Principles and Workflow
The Akida conversion [15] process follows a structured pipeline to ensure hardware compatibility and the preservation of model accuracy:
  • Operator Mapping: Each supported CNN layer (e.g., convolution, batch normalization, separable convolution, etc) is mapped to its spiking equivalent. For instance, ReLU activations are replaced by thresholding mechanisms that convert analog outputs to binary spike events. Layers incompatible with SNN operation (such as global average pooling or flatten) are substituted with spike-compatible alternatives, such as local pooling or spike-generating readout heads.
  • Quantization: All network weights and activations are quantized to a limited bit-width (8-bit), matching the precision constraints of the neuromorphic hardware. The quantization parameters are derived during the training and conversion phases to minimize information loss.
  • Temporal Spike Encoding: Continuous-valued activations from the CNN are converted into spike trains through rate coding or threshold-based event generation. The spike generation logic ensures that temporal information is preserved and the event-driven computation paradigm of SNNs is fully leveraged.
  • Resource and Constraint Adaptation: The MetaTF converter analyzes the input model to partition layers and neurons across the available hardware resources (neural cores, memory blocks) of the Akida chip, optimizing for parallelism, latency, and energy efficiency.
III-C2 Integration with Custom Backbone
Our proposed Quantization-Aware Network Architecture is specifically designed for seamless conversion. All intermediate activations are explicitly bounded and quantized, with network modules implemented in a form directly supported by Akida’s conversion pipeline. This ensures that no critical feature transformation is lost and that the functional mapping from input images to class predictions remains consistent between the CNN and its SNN counterpart.
III-C3 Conversion Output and Verification
Upon completion, the converter produces a deployable SNN model in Akida format. Model equivalence is empirically verified by comparing the output distributions of the original CNN and the converted SNN on a validation set. Any observed accuracy drop is mitigated through fine-tuning or incremental retraining on the SNN hardware, exploiting Akida’s support for on-chip learning.
III-D SNN Deployment, Optimization and Inference
The converted SNN model is deployed directly onto the BrainChip Akida neuromorphic processor for hardware-based inference. The deployment process comprises several steps: loading the SNN model into the hardware runtime environment, configuring input/output data streams, and initializing internal buffers and neuron state registers.
III-D1 Inference and Output Processing
Inference is performed in an event-driven manner, with input images encoded into spike trains and propagated through the SNN in a fully parallel fashion. For each test or validation sample, the model outputs spike counts or firing rates at the final output neurons, corresponding to the target classes. To robustly map temporal spike responses to class probabilities, we aggregate the spike counts $S_{c}(t)$ for each class $c$ within an integration window $T$, and apply a soft decision normalization:
$\hat{p}_{c}=\frac{\exp\left(\alpha\,\sum_{t=1}^{T}w_{t}S_{c}(t)\right)}{\sum_{k=1}^{C}\exp\left(\alpha\,\sum_{t=1}^{T}w_{t}S_{k}(t)\right)}$ (10)
where $w_{t}$ is an optional temporal weighting factor (e.g., $w_{t}=\exp(-\beta(T-t))$ for decaying integration), $\alpha$ is a scaling parameter, and $C$ is the number of output classes. Class prediction is then made as $\arg\max_{c}\hat{p}_{c}$.
III-D2 Parameter Calibration and On-Chip Optimization
After initial deployment, key runtime parameters—such as output spike thresholds, integration windows, and temporal pooling parameters—are empirically calibrated on a held-out validation set. To further optimize the class assignment and suppress spurious events, a threshold adaptation can be formulated as:
$\theta_{c}^{*}=\arg\min_{\theta}\left\{\sum_{i=1}^{N}\mathbb{I}\Big{[}\,y_{i}\neq\mathbb{I}\big{(}S_{c}^{(i)}>\theta\,\big{)}\Big{]}\right\}$ (11)
where $S_{c}^{(i)}$ is the total spike count for class $c$ on sample $i$, $y_{i}$ is the true label, and $\mathbb{I}[\cdot]$ is the indicator function. This allows data-driven threshold selection for each class. If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts.
IV Experiment
This section presents the datasets, experimental setup, and evaluation protocols, followed by detailed quantitative analyses of classification performance, ablation studies, and efficiency metrics under various deployment conditions.
IV-A Datasets
We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from Hospital Sultanah Bahiyah, Malaysia. Both sets reflect a range of lesion types and clinical diversity.
IV-A1 HAM10000 Dataset
The HAM10000 dataset [42] consists of 10,015 dermatoscopic RGB images labeled by expert dermatologists into seven categories: melanocytic nevus, melanoma, benign keratosis-like lesions, basal cell carcinoma, actinic keratosis/intraepithelial carcinoma, vascular lesions, and dermatofibroma. The images originate from diverse sources and exhibit significant class imbalance. For all experiments, we adopted a standard 70%/10%/20% train/validation/test split.
IV-A2 Hospital Sultanah Bahiyah Clinical Dataset
A proprietary clinical dataset was established in partnership with Hospital Sultanah Bahiyah, comprising 3,162 dermatoscopic images from 1,235 patients collected between June 2022 and February 2024. To ensure comparability with benchmark studies, lesion categories were selected to match the seven classes in HAM10000. Rare and unclassified lesions were excluded to maintain consistency in diagnostic labeling and facilitate joint evaluation. Each case was independently annotated by at least two board-certified dermatologists and histopathologically confirmed where possible. All images and metadata were anonymized following institutional ethical guidelines (approved protocol: NJG-2022-3233-CN). Dataset partitioning followed the same 70%/10%/20% train/validation/test split as HAM10000, stratified by disease category. This dataset provides a clinically diverse, real-world validation source for model generalization and robustness.
IV-B Experimental Setup
All model training, validation, and CNN-to-SNN conversion were conducted on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and NVIDIA RTX 3090 GPU, running Ubuntu 22.04 LTS. The software environment included Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1. Neuromorphic inference was performed on a BrainChip Akida AKD1000 PCIe board installed in the same system, with deployment and testing managed via the Akida Python API and default board settings.
IV-C Analysis of Classification Results on HAM10000 Dataset
Table I reports the precision, recall, and accuracy of the proposed model for each class on the HAM10000 test set. The model achieves consistent performance across all lesion categories, including minority classes. Notably, the architecture maintains stable results under limited training data and on-chip incremental learning, enabling effective clinical adaptation in evolving or low-resource scenarios. With a Top-1 accuracy of 91.6% and macro F1 of 82.4%, the system demonstrates effective discrimination of both common and rare lesions.
TABLE I: Class-wise precision, recall, F1 score, and accuracy of QANA on the HAM10000 test set.
Class Precision Recall F1 Accuracy
Actinic keratoses 0.890 0.933 0.911 0.933
Basal cell carcinoma 0.890 0.901 0.896 0.901
Benign keratosis-like lesions 0.866 0.853 0.859 0.853
Dermatofibroma 0.925 0.976 0.950 0.976
Melanocytic nevi 0.887 0.817 0.851 0.817
Vascular lesions 0.949 0.966 0.957 0.966
Melanoma 0.956 0.933 0.944 0.933
Average 0.909 0.911 0.910 0.910
TABLE II: Performance Comparison of Converted SNN Models on HAM10000
Model (SNN, Akida) Top-1 Accuracy (%) Macro F1 (%)
ResNet-50 [9] 85.7 76.4
DenseNet-121 [32] 86.5 77.2
Inception-v4 [33] 85.9 76.9
EfficientNet-B4 [43] 87.3 78.1
MobileNet-v2 [30] 83.4 74.7
SENet-154 [28] 86.9 77.8
Xception [11] 85.5 76.2
Multi-Scale Attention [35] 87.0 78.0
CNN Ensemble [44] 88.1 78.9
AKIDANet [15] 83.2 73.6
Ours 91.6 82.4
IV-D Classification Performance on HAM10000
To comprehensively evaluate the effectiveness of our neuromorphic skin lesion classification system, we conducted a series of controlled experiments in which a selection of state-of-the-art convolutional neural network (CNN) architectures were converted to spiking neural networks (SNNs) using our quantization-aware pipeline and executed on the same Akida hardware platform. This strategy ensures a fair comparison of all approaches under identical hardware constraints and SNN deployment settings. The evaluated models include both canonical CNN baselines and advanced architectures commonly used in medical image analysis.
As shown in Table II, all evaluated CNN models show a decrease in accuracy and macro F1 score after conversion to SNN and deployment on neuromorphic hardware, with Top-1 accuracy values ranging from 83.2% to 88.1%. In contrast, our proposed model achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%, outperforming all baselines under the same SNN deployment conditions. This improvement demonstrates the benefit of our quantization-aware architecture and optimized network design for event-driven inference. The results confirm the practical utility of our method for real-time, resource-constrained medical applications, supporting efficient and accurate classification in portable diagnostic systems.
TABLE III: Per-image inference latency and energy consumption of all models on the HAM10000 test set. CNN baselines are measured on an NVIDIA RTX 3090 (GPU) and Intel Xeon Gold 6226R (CPU); SNNs are measured on BrainChip Akida AKD1000. All values are averaged over 10,000 images. The rightmost columns show the percentage reduction achieved by SNNs on Akida compared to the corresponding CNN (GPU) version.
Model CNN (GPU) CNN (CPU) SNN (Akida) Relative Reduction(%)
Latency (ms) Energy (mJ) Latency (ms) Energy (mJ) Latency (ms) Energy (mJ) Latency Energy
ResNet-50 [9] 12.1 175.2 57.9 923.3 2.8 3.3 76.9 98.1
DenseNet-121 [32] 14.7 199.6 68.4 1075.2 3.1 3.5 78.9 98.2
Inception-v4 [33] 16.8 218.5 82.1 1237.5 3.5 4.1 79.2 98.1
EfficientNet-B4 [43] 18.9 242.1 93.6 1345.6 4.0 4.6 78.8 98.1
MobileNet-v2 [30] 6.9 97.5 68.2 1082.1 2.2 2.6 68.1 97.3
SENet-154 [28] 17.3 236.8 89.4 1312.4 3.8 4.4 78.0 98.1
Xception [11] 10.7 151.2 51.5 863.7 2.6 3.1 75.7 97.9
Multi-Scale Attention [35] 19.8 251.7 97.7 1391.5 4.2 5.0 78.8 98.0
CNN Ensemble [44] 36.1 450.5 157.2 2177.1 7.5 8.6 79.2 98.1
QANA (Ours) 27.6 163.1 83.9 841.5 1.5 1.7 94.6 98.6
TABLE IV: Ablation study of core modules in our model on the HAM10000 test set. Metrics are reported as percentages (%). Each row shows the incremental addition of modules.
Configuration Ghost Block ECA SE Quant. Head SMOTE Inc. Learn Accuracy Recall Precision F1 Score AUC-ROC
Baseline 74.1 71.4 71.9 71.6 77.3
+ Ghost Block 72.3 70.2 70.0 70.6 70.9
+ ECA 88.7 85.8 87.2 86.5 90.7
+ SE 89.8 87.7 88.1 87.8 91.5
+ Augmentation 90.4 88.1 89.1 88.6 92.1
+ SMOTE 91.0 89.2 90.0 89.6 92.7
+ Incremental Learning 91.6 90.7 91.2 90.9 93.4
IV-E Inference Speed and Energy Consumption
We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures. All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA-RTX-3090-GPU and an Intel Xeon CPU. For each model, the reported inference latency corresponds to the average per-image processing time over 10,000 test samples. Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida. Table III presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations. Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures. When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 99.0% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
TABLE V: Performance comparison of converted SNN models on the Clinical Dataset.
Model (SNN, Akida) Top-1 Accuracy (%) Macro F1 (%)
ResNet-50 [9] 84.6 75.3
DenseNet-121 [32] 85.7 76.2
Inception-v4 [33] 85.2 75.8
EfficientNet-B4 [43] 86.3 77.0
MobileNet-v2 [30] 82.8 73.7
SENet-154 [28] 85.4 76.6
Xception [11] 84.2 74.8
Multi-Scale Attention [35] 86.5 77.2
CNN Ensemble [44] 87.6 78.1
AKIDANet [15] 81.9 71.5
Ours 90.8 81.7
IV-F Ablation Study of Model Components
Table IV reports the ablation results on the HAM10000 test set, illustrating the contribution of each core module to the overall performance. Modules were incrementally enabled to measure their isolated and cumulative effects on classification metrics, including accuracy, recall, precision, F1 score, and AUC-ROC (%). The results demonstrate that each module provides consistent improvements, with the complete model achieving the highest accuracy and F1 score.
IV-G Classification Performance on Clinical Dataset
We evaluated the proposed neuromorphic skin lesion classification system and several representative CNN architectures, all converted to SNNs and deployed on the Akida hardware platform. Table V presents the Top-1 accuracy and macro F1 score for each model under the same deployment conditions. Our model achieves the highest accuracy and macro F1 score among all tested methods, confirming its effectiveness and robustness for neuromorphic inference in clinical scenarios.
V Conclusion
In this paper, we proposed QANA, a quantization-aware neuromorphic framework for skin lesion classification on edge devices. Extensive experiments on the large-scale HAM10000 benchmark and a real-world clinical dataset show that QANA achieves state-of-the-art accuracy (91.6% Top-1, 82.4% macro F1 on HAM10000; 90.8%/81.7% on the clinical set) while enabling real-time and energy-efficient inference on the BrainChip Akida platform (1.5 ms latency, 1.7 mJ per image). These results demonstrate that QANA is highly effective for portable medical analysis and AI deployment in dermatology under limited computing resources.
VI Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62172336 and 62032018). The authors gratefully acknowledge BrainChip Holdings Ltd. for providing technical support and the Akida AKD1000 hardware platform, whose powerful neuromorphic computing capabilities enabled strong performance of the SNN model. The authors also extend their appreciation to Dr. Atif Mansoor, Dr. Bo Miao and their teams for their preliminary contributions to this research.

### v2

@@
 Abstract
-Accurate and efficient skin lesion classification on edge devices is critical for accessible dermatological care but remains challenging due to computational, energy, and privacy constraints.
-We introduce QANA, a novel quantization-aware neuromorphic architecture for incremental skin lesion classification on resource-limited hardware.
+On-device skin lesion edge classification remains hindered by the lack of efficient architectures compatible with neuromorphic deployment, as conventional CNN-to-SNN conversion suffers from structural incompatibility, performance degradation on rare diseases, and absence of continual learning.
+To address these limitations, we propose an end-to-end pipeline that incorporates a quantization-aware CNN backbone explicitly designed for conversion-preserving mapping to SNNs and neuromorphic deployment, culminating in the QANA architecture optimized for low-latency, energy-efficient, and incremental inference.
 QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference.
@@
 These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
-Index Terms:
-Resource-Constrained Devices, Edge Computing, Neuromorphic Computing, Low Latency, Energy Efficiency.
 I Introduction
@@
 In addition, converted SNNs frequently experience significant accuracy loss on small, imbalanced medical datasets due to quantization effects and limited capacity for capturing subtle pathological features [23].
-In this work, we present an end-to-end pipeline for efficient and incremental skin lesion classification on neuromorphic hardware.
-As illustrated in Fig.
-1, our approach consists of four main stages: (1) data preprocessing, including image quality screening, augmentation, and SMOTE-based class balancing; (2) construction of a quantization-aware neural network with stacked Ghost modules, channel attention, and a spike-compatible output head; (3) automated conversion of the trained network to a spiking neural network (SNN) using the Akida MetaTF toolkit; and (4) deployment and optimization on the BrainChip Akida neuromorphic platform for low-latency, energy-efficient inference.
-The core QANA architecture is specifically designed for hardware compatibility, enabling direct and lossless translation to spike-based event-driven computation.
-This pipeline directly addresses the major challenges of edge deployment, including computational constraints, class imbalance, limited data, and the need for incremental model updates.
-Experimental results on both the HAM10000 public benchmark and a real-world clinical dataset demonstrate the practical advantages of our approach.
-QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% / 81.7% on the clinical dataset.
-On Akida hardware, the system delivers 1.5 ms inference latency and 1.7 mJ energy consumption per image, surpassing state-of-the-art CNN-to-SNN conversion baselines in both accuracy and efficiency.
-These findings confirm the effectiveness of QANA for resource-constrained, real-time medical image analysis.
-II Related Work
-Prior studies on neuromorphic hardware and Spiking Neural Networks (SNNs) have highlighted their suitability for edge inference and continual learning in constrained environments [14].
-Platforms such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17] demonstrate ultra-low-power operation and support on-chip incremental learning, enabling model updates without full retraining.
-These features are advantageous for medical imaging tasks on edge devices.
-However, existing works primarily focus on general benchmarks like MNIST [24] or CIFAR [25] and rarely address domain-specific challenges in dermatology [26].
-Furthermore, in medical image contexts, edge Spiking Neural learning methods often suffer from catastrophic forgetting and fail to preserve performance on previously learned classes when new samples are introduced [27].
-CNN-to-SNN conversion has become a common strategy to leverage mature CNN architectures—ResNet, DenseNet—for deployment on neuromorphic processors [28].
-Conversion toolkits, including Akida’s CNN2SNN [29], deliver compatibility by quantizing weights and replacing activation functions with spike-based equivalents.
-While some studies report preserved accuracy on standard datasets [30], performance significantly degrades on small or imbalanced medical datasets due to batching overfitting, quantization noise, and inability to capture rare lesion features [15].
-Limitations in mapping CNN operations—such as batch normalization, global average pooling, and multi-bit activations—to spiking neurons further compound accuracy loss during conversion [31].
+We propose a quantization-aware neuromorphic architecture (QANA) within an end-to-end pipeline (Fig.
+1) to overcome these challenge.
+QANA comprises three essential modules: (1) iterative feature extraction and downsampling block utilizing stacked Ghost modules, spatially-aware efficient channel attention (SA-ECA), and residual connections to reduce computation and inference latency, thereby directly facilitating effective CNN-to-SNN conversion; (2) spike-compatible feature transformation block employing quantization-aware normalization, bounded activations, and squeeze-and-excitation (SE) blocks, ensuring hardware compatibility and efficient incremental learning; and (3) quantized output projection block explicitly optimized for conversion-preserving CNN-to-SNN mapping and low-power neuromorphic deployment.
+These components collectively address key conversion limitations, preserve high accuracy for rare diseases, and efficiently support incremental model updates on edge devices.
+Evaluations on the HAM10000 benchmark dataset show that QANA achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%.
+Further validation on a clinical dataset confirms robustness, achieving a Top-1 accuracy of 90.8% and a macro F1 score of 81.7%.
+When deployed on the Akida neuromorphic hardware platform, QANA attains an inference latency of 1.5 ms per image and energy consumption of 1.7 mJ, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, surpassing state-of-the-art CNN-to-SNN conversion methods in computational efficiency.
+These outcomes verify QANA’s practical suitability for resource-constrained and privacy-sensitive medical diagnostic applications.
+Our main contributions are:
+  • Conversion-preserving architecture.
+We design QANA with bounded activations, separable and ghost convolutions, and lightweight channel attention with SE gating, ensuring all operators lower one-to-one to 8-bit spiking kernels.
+  • Deployment-aligned training and conversion.
+We provide a graph-preserving flow with batch-norm fusion and affine 8-bit quantization, combined with QAT, embedding-space SMOTE with safety filtering, and a lightweight on-chip incremental readout update, all under a fixed $64{\times}64$ input and matched preprocessing.
+  • Edge evaluation.
+On Akida and under identical preprocessing and quantization for baselines, QANA achieves 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical set, with 1.5 ms latency and 1.7 mJ per image.
 Figure 2: Detailed architecture of our QANA, which performs iterative feature extraction using stacked Ghost modules, ECA, and residual blocks, followed by spike-compatible transformation with batch normalization, ReLU activation, and Squeeze-and-Excitation (SE) block.
 The output is then quantized and projected to class logits for SNN deployment.
-In parallel, conventional CNN-based methods for skin lesion classification, such as those using DenseNet-121 [32] or Inception-v4 [33], show high performance on benchmark datasets like ISIC.
-Despite their effectiveness, these models necessitate continuous retraining on cloud infrastructure to address domain shifts, new lesion types, or demographic variations [18].
-Such retraining incurs latency, high computational cost, and data privacy concerns [34].
-Moreover, general CNN models often fail to generalize to underrepresented conditions [35], such as rare tumors or images from diverse skin tones, highlighting the need for adaptive and privacy-preserving on-device solutions.
-These limitations motivate the development of neuromorphic frameworks that maintain diagnostic performance while supporting incremental learning and efficient operation in resource-constrained settings.
+II Related Work
+This section surveys the landscape for edge dermatology.
+We first outline neuromorphic hardware capabilities and constraints that govern feasible model design.
+We then compare ANN-to-SNN conversion with direct SNN training, including spike encoding and sparsity control, and finally review CNN-based dermatology systems as accuracy baselines.
+Across these threads we emphasize limitations under small and imbalanced datasets and strict latency and energy budgets, which together define the problem setting.
+II-A Neuromorphic Hardware for Edge Imaging
+Neuromorphic processors execute event-driven computation with low power and support on-chip adaptation, which matches the requirements of on-device medical imaging [14].
+Akida, TrueNorth, and Loihi provide hardware support for spiking inference and incremental learning [15, 16, 17].
+At the same time, these platforms constrain model design through limited operator sets, fixed-point quantization, and resource budgets that restrict input resolution and intermediate state.
+Related methods must therefore account for quantization and operator compatibility during both training and deployment.
+II-B ANN-to-SNN Conversion
+A prevalent strategy trains a conventional CNN and converts it to an SNN by mapping activations to spikes with calibrated thresholds and quantized ranges [24].
+Conversion enables reuse of mature backbones but introduces several sources of error.
+Batch normalization and global pooling do not admit direct spike-based equivalents, which forces structural edits and alters feature statistics.
+Rate-coded readouts require temporal integration that increases latency and makes predictions sensitive to threshold and timestep calibration, which is difficult to tune on small and imbalanced datasets.
+Empirical studies report measurable accuracy drops after conversion under these conditions, indicating that quantization and operator mismatch are primary failure modes on edge workloads [15].
+II-C Directly Trained SNNs and Spike Encoding
+Direct training replaces the non-differentiable spike with surrogate derivatives and optimizes spiking dynamics end-to-end.
+SLAYER derives a time-domain error assignment rule for convolutional SNNs with learnable delays [25], and surrogate-gradient frameworks formalize stable optimization for deep networks [26].
+Subsequent work improves few-timestep accuracy with residual spiking blocks [27] and temporal efficient training that couples the loss across discrete integration steps [28].
+Despite these advances, training remains sensitive to surrogate choice, thresholds, and timestep schedules.
+Performance depends on spike encoding.
+Time-to-first-spike and related temporal codes reduce spikes and latency [29, 30], yet they complicate calibration on hardware designed for rate accumulation and may reduce robustness on fine-grained classes when data are scarce.
+Activity regularization controls event rates and energy [31, 25], but aggressive sparsity can underfit rare patterns.
+These factors make it difficult to guarantee accuracy, latency, and energy simultaneously under fixed hardware constraints.
+II-D CNN-Based Dermatology Models
+Dermatology classifiers based on DenseNet-121 and Inception-v4 reach strong results on ISIC-style datasets [32, 33].
+In practice, deployment often depends on cloud retraining to handle distribution shift and rare categories, which introduces latency and privacy risk [18, 34].
+Training data are class-imbalanced and images vary across devices and skin tones, which limits generalization to underrepresented conditions [35].
+These limitations indicate that comparative evaluation on edge hardware should consider both conversion-based pipelines and directly trained SNNs with temporal encodings and sparsity control, so that the impact of quantization, operator compatibility, and encoding choices on accuracy, latency, and energy can be isolated under identical deployment settings.
 III Methodology
@@
 III-A1 Image Preprocessing
-All dermatoscopic images are first screened for quality.
-Corrupted, low-resolution, or artifact-contaminated images are excluded using automated checks based on image metadata and pixel statistics, followed by manual review when ambiguity remains.
-For compatibility with the Akida neuromorphic platform, all accepted images are resized to $64\times 64$ pixels using bilinear interpolation, which is the maximum supported resolution under current on-chip memory constraints.
-Pixel intensities are normalized channel-wise to the range $[0,1]$ to ensure consistent input distribution.
-No denoising or inpainting operations are performed, as quality control steps eliminate samples with significant artifacts.
+All dermatoscopic images undergo automated integrity checks followed by a brief manual review when needed.
+Inputs are processed at $64\times 64$ to match the Akida on-chip memory budget used in our deployment.
+To retain lesion content at this resolution, we apply Grad-CAM on the pretrained backbone to obtain a coarse saliency map.
+A tight square crop is centered on the top-activated region and resized with antialiased area resampling to $64\times 64$.
+If the saliency peak lies within 10% of the image center or the saliency map is flat (coefficient of variation <0.05<0.05), we fall back to a centered square crop.
+Pixel intensities are normalized per channel to $[0,1]$.
+No denoising or inpainting is used.
 III-A2 Data Augmentation
-To mitigate overfitting and increase the diversity of the training dataset, each image in the training split undergoes stochastic augmentation with the following operations: brightness adjustment (random factor sampled uniformly from $[0.7,1.3]$), contrast modification ($[0.8,1.2]$), random horizontal and vertical flipping (each with probability 0.5), hue shift (random shift in $[-0.08,0.08]$ in HSV space), and saturation variation ($[0.85,1.15]$).
-Each augmentation is applied independently to each sample with the specified probability.
-All augmentation parameters are determined by fixed random seeds for reproducibility.
-These transformations simulate real-world imaging variation while preserving lesion semantics.
+Augmentations preserve dermoscopic color cues.
+We use random horizontal and vertical flips (probability 0.5).
+Luminance is perturbed on the CIELAB $L$ channel with a multiplicative factor in $[0.95,1.05]$ or a gamma factor in $[0.97,1.03]$.
+Contrast is adjusted in $[0.95,1.05]$.
+No hue shift is applied.
+A gray-world normalization fitted on the training split is used during training only.
+All random draws use fixed seeds for reproducibility.
 III-A3 Synthetic Minority Oversampling (SMOTE)
-To address class imbalance, especially in rare lesion categories, the Synthetic Minority Oversampling Technique (SMOTE) [36] is applied to the training data.
-Given a minority class dataset $\mathcal{S}=\{x_{1},x_{2},...,x_{n}\}$ in $\mathbb{R}^{d}$, for each $x_{i}\in\mathcal{S}$, the $k$ nearest neighbors $\mathcal{N}_{i}=\{x_{i_{1}},...,x_{i_{k}}\}$ are identified in feature space via Euclidean distance.
-For each synthetic sample, a random neighbor $x_{i_{j}}$ is selected, and a new sample is generated by:
-$x_{\text{new}}=x_{i}+\lambda\cdot(x_{i_{j}}-x_{i}),\quad\lambda\sim\mathcal{U}(0,1)$ (1)
-where $\lambda$ is independently sampled for each feature dimension. Repeating this process for all $x_{i}$ and multiple neighbors results in the augmented set:
-$\mathcal{S}_{\text{aug}}=\mathcal{S}\cup\left\{x_{i}+\lambda^{(m)}\cdot(x_{i_{j_{m}}}-x_{i})\mid x_{i}\in\mathcal{S},\,1\leq m\leq M\right\}$ (2)
-where $M$ is the number of synthetic samples generated per original sample.
-All images are maintained at $64\times 64$ resolution to conform with Akida’s memory requirements throughout oversampling.
-SMOTE is applied only to the training set to preserve test set integrity and eliminate information leakage.
+We perform SMOTE strictly in the learned embedding space rather than on raw pixels.
+For each training image $x_{i}$ from class $c$, the backbone and spike-compatible transform produce a penultimate representation $\mathbf{r}_{i}\in\mathbb{R}^{4096}$ after global flattening.
+Let $\mathcal{I}_{c}$ be the index set of class $c$ and $\{\mathbf{r}_{i}\}_{i\in\mathcal{I}_{c}}$ the corresponding embeddings.
+Before neighbor search, embeddings are standardized on the training split only to avoid information leakage.
+A PCA whitening is then fitted on the same split and applied to stabilize distances in high dimension.
+Writing the sample covariance as $\mathbf{\Sigma}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ and retaining the top $d=256$ components, the whitened vector is
+$\widetilde{\mathbf{r}}_{i}=\mathbf{\Lambda}_{d}^{-\frac{1}{2}}\mathbf{U}_{d}^{\top}(\mathbf{r}_{i}-\boldsymbol{\mu}),$ (1)
+where $\boldsymbol{\mu}$ is the training mean, $\mathbf{U}_{d}\in\mathbb{R}^{4096\times d}$ the leading eigenvectors and $\mathbf{\Lambda}_{d}\in\mathbb{R}^{d\times d}$ their eigenvalues.
+Nearest neighbors are computed with Euclidean distance in the whitened space; for each $i\in\mathcal{I}_{c}$ we form a $k$-NN set $\mathcal{N}_{i}$ with $k=5$.
+To reduce the influence of outliers, we keep only “safe” minority anchors whose local median neighbor distance is below the class-specific 90th percentile:
+$\mathcal{S}_{c}^{\text{safe}}=\Big\{\,i\in\mathcal{I}_{c}\ \Big|$ $\operatorname{med}_{j\in\mathcal{N}_{i}}\big\lVert\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}$ (2)
+$\leq Q_{0.9}\!\left(\big\{\operatorname{med}_{j\in\mathcal{N}_{u}}\big\lVert\widetilde{\mathbf{r}}_{u}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}\big\}_{u\in\mathcal{I}_{c}}\right)\Big\}.$
+For each minority class $c$ with $N_{c}=|\mathcal{I}_{c}|$ and $N_{\max}=\max_{c^{\prime}}N_{c^{\prime}}$, we synthesize to a target level $\tau N_{\max}$ with $\tau=0.85$ to avoid distribution collapse.
+The number of synthetic embeddings is
+$M_{c}=\max\!\big(0,\;\lfloor\tau N_{\max}-N_{c}\rfloor\big).$ (3)
+Each synthetic vector is generated from an anchor $i\in\mathcal{S}_{c}^{\text{safe}}$ and a random neighbor $j\in\mathcal{N}_{i}$ by a convex interpolation and a small, locally aligned perturbation to improve diversity.
+Let $\mathbf{C}_{i}$ be the empirical covariance of $\{\widetilde{\mathbf{r}}_{u}\}_{u\in\{i\}\cup\mathcal{N}_{i}}$, and let $\mathbf{U}_{i}^{(m)}$ contain its $m=8$ principal directions.
+With $\delta_{i}=\operatorname{med}_{j\in\mathcal{N}_{i}}\|\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\|_{2}$ and $\sigma_{c}=0.05\,\operatorname{med}_{u\in\mathcal{I}_{c}}\delta_{u}$, we sample
+$\widetilde{\mathbf{r}}_{\text{new}}=\widetilde{\mathbf{r}}_{i}+\eta\big(\widetilde{\mathbf{r}}_{j}-\widetilde{\mathbf{r}}_{i}\big)+\sigma_{c}\,\mathbf{U}_{i}^{(m)}\boldsymbol{\xi},\quad\eta\sim\mathcal{U}(0,1),\;\boldsymbol{\xi}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m}).$ (4)
+The synthetic embedding is then mapped back to the original feature space by the inverse whitening:
+$\mathbf{r}_{\text{new}}=\boldsymbol{\mu}+\mathbf{U}_{d}\mathbf{\Lambda}_{d}^{\frac{1}{2}}\widetilde{\mathbf{r}}_{\text{new}}.$ (5)
+Synthetic data are not images.
+They are injected at the penultimate layer as additional feature vectors and used to train the quantized output projection with cross-entropy.
+During this stage the backbone up to the spike-compatible transform is frozen; only the classifier head is updated for 5 epochs with mixed real and synthetic embeddings.
+A subsequent short fine-tuning on real data only re-enables the last block and the head to reconcile the feature distribution.
+Input images are still resized to $64\times 64$ prior to feature extraction for hardware constraints, but no pixel-space mixing or image synthesis is performed at any time.
+Figure 3: Schematic of the Ghost module.
+The input feature map is first processed by a lightweight convolution to extract a reduced set of primary features with channel size $\mu C$, where $C$ is the target output dimensionality and $\mu\in(0,1)$ is a tunable ratio.
+Subsequently, inexpensive operations are applied to the primary features to generate additional ghost features of size $(1-\mu)C$.
+These are concatenated along the channel axis to form the final output of size $C$.
 III-B Quantization-Aware Network Architecture
@@
 This quantization-aware step is crucial for downstream SNN compatibility and robust generalization.
-Figure 3: Schematic of the Ghost module.
-The input feature map is first processed by a lightweight convolution to extract a reduced set of primary features with channel size $\mu C$, where $C$ is the target output dimensionality and $\mu\in(0,1)$ is a tunable ratio.
-Subsequently, inexpensive operations are applied to the primary features to generate additional ghost features of size $(1-\mu)C$.
-These are concatenated along the channel axis to form the final output of size $C$.
 To capture cross-channel correlations, each block integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) mechanism, inspired by Efficient Channel Attention (ECA) mechanism [37].
@@
 The resulting attention weights are used to rescale the input feature channels, enhancing discriminative information with minimal computational overhead.
-Figure 5: Illustration of the Squeeze-and-Excitation (SE) block.
-The input feature map undergoes global pooling, followed by two fully connected layers with ReLU and sigmoid activations to compute channel-wise weights.
-The original feature map is then rescaled by these weights, enabling adaptive recalibration of channel responses.
 III-B2 Spike-Compatible Feature Transformation
@@
 A SeparableConv2D $(3\times 3,256)$ [38] generates higher-dimensional features, followed by quantization-aware normalization and bounded activation:
-$\widehat{F}=\min\left(1,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (7)
+$\widehat{F}=\min\left(6,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (10)
 This mapping guarantees that all activations lie in $[0,1]$, which is both compatible with spike encoding and preserves information for subsequent inference.
@@
 Each feature channel is then scaled by $s_{c}$, which adaptively modulates discriminative capacity and provides additional regularization for small, imbalanced datasets.
+Figure 5: Illustration of the Squeeze-and-Excitation (SE) block.
+The input feature map undergoes global pooling, followed by two fully connected layers with ReLU and sigmoid activations to compute channel-wise weights.
+The original feature map is then rescaled by these weights, enabling adaptive recalibration of channel responses.
 III-B3 Quantized Output Projection
@@
 III-C CNN-to-SNN Conversion for Neuromorphic Deployment
-The conversion of a trained convolutional neural network (CNN) into a spiking neural network (SNN) is a critical step for enabling event-driven inference on neuromorphic hardware.
-In this work, we utilize the Akida MetaTF toolkit [41, 15] to perform an automated and quantization-aware transformation of the CNN backbone described above into an SNN model suitable for direct deployment on the BrainChip Akida processor.
-III-C1 Conversion Principles and Workflow
-The Akida conversion [15] process follows a structured pipeline to ensure hardware compatibility and the preservation of model accuracy:
-  • Operator Mapping: Each supported CNN layer (e.g., convolution, batch normalization, separable convolution, etc) is mapped to its spiking equivalent.
-For instance, ReLU activations are replaced by thresholding mechanisms that convert analog outputs to binary spike events.
-Layers incompatible with SNN operation (such as global average pooling or flatten) are substituted with spike-compatible alternatives, such as local pooling or spike-generating readout heads.
-  • Quantization: All network weights and activations are quantized to a limited bit-width (8-bit), matching the precision constraints of the neuromorphic hardware.
-The quantization parameters are derived during the training and conversion phases to minimize information loss.
-  • Temporal Spike Encoding: Continuous-valued activations from the CNN are converted into spike trains through rate coding or threshold-based event generation.
-The spike generation logic ensures that temporal information is preserved and the event-driven computation paradigm of SNNs is fully leveraged.
-  • Resource and Constraint Adaptation: The MetaTF converter analyzes the input model to partition layers and neurons across the available hardware resources (neural cores, memory blocks) of the Akida chip, optimizing for parallelism, latency, and energy efficiency.
-III-C2 Integration with Custom Backbone
-Our proposed Quantization-Aware Network Architecture is specifically designed for seamless conversion.
-All intermediate activations are explicitly bounded and quantized, with network modules implemented in a form directly supported by Akida’s conversion pipeline.
-This ensures that no critical feature transformation is lost and that the functional mapping from input images to class predictions remains consistent between the CNN and its SNN counterpart.
-III-C3 Conversion Output and Verification
-Upon completion, the converter produces a deployable SNN model in Akida format.
-Model equivalence is empirically verified by comparing the output distributions of the original CNN and the converted SNN on a validation set.
-Any observed accuracy drop is mitigated through fine-tuning or incremental retraining on the SNN hardware, exploiting Akida’s support for on-chip learning.
+We convert the trained quantization-aware CNN to an event-driven SNN using a graph-preserving flow with three stages: static reparameterization, integer quantization with calibration or QAT, and hardware-constrained code generation.
+First, all Batch Normalization layers are eliminated by fusing their affine statistics into the preceding convolution or pointwise projection.
+For a convolution output channel $c$ with kernel $\mathbf{W}_{c}$ and bias $b_{c}$ and BN parameters $(\gamma_{c},\beta_{c},\mu_{c},\sigma_{c}^{2},\epsilon)$, the fused weights and bias are
+$\mathbf{W}_{c}^{\prime}=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\epsilon}}\mathbf{W}_{c},\qquad b_{c}^{\prime}=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\epsilon}}\left(b_{c}-\mu_{c}\right)+\beta_{c}.$ (13)
+Depthwise and separable convolutions are treated identically channel-wise.
+Nonlinearities not in the target activation set are replaced during training by bounded ReLU6 so that all post-activation tensors are in $[0,6]$, which simplifies downstream spike generation and avoids undefined mappings in the spiking domain [40, 24].
+Global average pooling is lowered to an integer spatial reduction with compile-time division embedded in the output scale.
+Flatten is treated as a metadata reshape and does not generate compute kernels.
+Softmax is applied on the host side after spike integration[24].
+Quantization follows an affine, per-tensor, uniform scheme for both weights and activations with 8-bit fixed-point.
+Let $a$ denote a real activation and $(s_{a},z_{a})$ the scale and zero-point obtained by calibration on a held-out set or learned with quantization-aware training.
+The integer activation and its error are
+$a_{q}=\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{a}{s_{a}}\right)+z_{a},\,0,\,255\right),e_{a}=a-\left(a_{q}-z_{a}\right)s_{a}.$ (14)
+During SNN emission, each neuron integrates the quantized drive and generates a binary event whenever the internal accumulator crosses a programmable threshold $\tau$; resetting is subtractive.
+With state $u_{t}$ and drive $a_{q}$ the discrete-time dynamics are
+$s_{t}=\mathbb{I}\!\left(u_{t-1}+a_{q}\geq\tau\right),\qquad u_{t}=\left(u_{t-1}+a_{q}\right)-\tau s_{t},$ (15)
+and rate readout uses the spike count within an integration window.
+For ReLU6, $s_{a}=255/6$ and $z_{a}=0$ make the activation-range utilization explicit and fix the time-to-first-spike under a given $\tau$ [24].
+Operator lowering is rule-based and deterministic.
+Convolution, depthwise convolution, separable convolution, average pooling, max pooling, elementwise clipping and integer linear layers are mapped one-to-one to supported spiking kernels.
+Global average pooling is lowered to an integer spatial reduction with compile-time division by the pooling area embedded in the output scale.
+Flatten is a metadata reshape that preserves memory layout and does not generate compute kernels.
+Unsupported nonlinearities such as Swish or hard-Swish are removed at training time and replaced by ReLU6.
+Squeeze-and-Excitation is realized by keeping the global pooling and two low-rank linear projections in the quantized graph and folding the final channel-wise gating into the subsequent pointwise projection when possible, otherwise emitting a per-channel integer rescale.
+Residual additions are integer tensor adds after scale alignment by rational multipliers that are absorbed into the successor scale.
+The converter rejects any subgraph that would require a floating-point dequantize.
+To make the comparison across backbones fair, a model is deemed convertible only if the compiled spiking graph contains no dequantize fallback and every tensor is quantized to 8-bit throughout the feature extraction and readout path; otherwise the model is excluded from neuromorphic evaluation and reported as a failed conversion.
+Fairness metrics and logs are computed on the same train and test splits and under identical preprocessing.
+We retrain each CNN baseline with the same input resolution, the same augmentation policy, the same learning-rate schedule, and the same 8-bit quantization configuration, then convert and deploy the resulting SNN.
+Conversion success rate is defined on the quantized graph $\mathcal{G}$ as
+$r_{\mathrm{succ}}=1-\frac{\lvert\mathcal{V}_{\mathrm{fp32}}\rvert}{\lvert\mathcal{V}\rvert}$, (16)
+where $\mathcal{V}$ is the node set and $\mathcal{V}_{\mathrm{fp32}}$ are nodes that would require dequantization and thus cannot be placed on the neuromorphic cores.
+Accuracy drop is reported as $\Delta\mathrm{Acc}=\mathrm{Acc}_{\mathrm{SNN}}-\mathrm{Acc}_{\mathrm{CNN}}$ and $\Delta\mathrm{F1}=\mathrm{F1}_{\mathrm{SNN}}-\mathrm{F1}_{\mathrm{CNN}}$ on the same test set.
+For transparency, we release per-model conversion manifests listing, for every layer, the integer scale and zero-point, the applied reparameterization, and the final kernel selected by the code generator.
+The mapping rules in (13)–(16) and the practice of bounded activations to guarantee spike-compatibility follow established CNN-to-SNN conversion principles and 8-bit QAT literature.
 III-D SNN Deployment, Optimization and Inference
@@
 If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts.
+For higher-detail cases, a $2\times 2$ tiling fallback is triggered when $\max_{c}\hat{p}_{c}<0.60$.
+Four overlapping $64\times 64$ patches (20% stride overlap) are evaluated and logits averaged.
+On HAM10000, the trigger rate is 18%.
 IV Experiment
@@
 IV-A Datasets
-We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from Hospital Sultanah Bahiyah, Malaysia. Both sets reflect a range of lesion types and clinical diversity.
+We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from local hospital. Both sets reflect a range of lesion types and clinical diversity.
 IV-A1 HAM10000 Dataset
@@
 For all experiments, we adopted a standard 70%/10%/20% train/validation/test split.
-IV-A2 Hospital Sultanah Bahiyah Clinical Dataset
-A proprietary clinical dataset was established in partnership with Hospital Sultanah Bahiyah, comprising 3,162 dermatoscopic images from 1,235 patients collected between June 2022 and February 2024.
-To ensure comparability with benchmark studies, lesion categories were selected to match the seven classes in HAM10000.
-Rare and unclassified lesions were excluded to maintain consistency in diagnostic labeling and facilitate joint evaluation.
-Each case was independently annotated by at least two board-certified dermatologists and histopathologically confirmed where possible.
-All images and metadata were anonymized following institutional ethical guidelines (approved protocol: NJG-2022-3233-CN).
-Dataset partitioning followed the same 70%/10%/20% train/validation/test split as HAM10000, stratified by disease category.
-This dataset provides a clinically diverse, real-world validation source for model generalization and robustness.
+IV-A2 local Clinical Dataset
+A non-public clinical dermatoscopic dataset was provided by a partner hospital under a data-use agreement.
+The dataset comprises 3,162 images from 1,835 patients acquired during routine care between June 2022 and February 2024.
+No human experiments or prospective data collection were conducted by the authors.
+Lesion categories were harmonized to the seven HAM10000 classes; rare or unclassified cases were excluded to maintain consistent labeling.
+Each case was independently annotated by at least two board-certified dermatologists, with histopathology confirmation when available.
+All records were de-identified before access and used solely for secondary analysis.
+Dataset partitioning followed the same $70\%/10\%/20\%$ train/validation/test split as HAM10000, stratified by disease category, providing a clinically diverse, real-world validation source for model generalization.
 IV-B Experimental Setup
-All model training, validation, and CNN-to-SNN conversion were conducted on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and NVIDIA RTX 3090 GPU, running Ubuntu 22.04 LTS.
-The software environment included Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1.
-Neuromorphic inference was performed on a BrainChip Akida AKD1000 PCIe board installed in the same system, with deployment and testing managed via the Akida Python API and default board settings.
+All training and conversion are performed on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and an NVIDIA RTX 3090, running Ubuntu 22.04 LTS.
+Software includes Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1.
+Unless otherwise stated, batch size is 1 for latency and energy measurement.
+We use AdamW (weight decay $10^{-4}$), initial learning rate $3\times 10^{-4}$ with cosine decay for 120 epochs, and early stopping on validation macro F1.
+Quantization-aware training uses 8-bit affine per-tensor quantization for weights and activations with ReLU6 clipping.
+Neuromorphic inference uses a BrainChip Akida AKD1000 PCIe board with default runtime settings.
+GPU energy is measured as the board power above idle (averaged by NVML at 20 Hz) multiplied by the per-image inference time; Akida energy is measured analogously by on-board power monitoring.
+Reported Akida 1.5 ms corresponds to on-chip execution time; host preprocessing and I/O add <0.2<0.2 ms on average and are reported separately when stated.
+TABLE I: Class-wise precision, recall, F1, and per-class accuracy on the HAM10000 test set. The macro F1 equals the unweighted mean of per-class F1.
+Class Precision Recall F1 Accuracy
+Actinic keratoses 0.840 0.820 0.830 0.820
+Basal cell carcinoma 0.830 0.820 0.820 0.820
+Benign keratosis-like 0.800 0.780 0.790 0.780
+Dermatofibroma 0.900 0.860 0.880 0.860
+Melanocytic nevi 0.800 0.740 0.770 0.740
+Vascular lesions 0.870 0.850 0.860 0.850
+Melanoma 0.830 0.810 0.820 0.810
+Macro average 0.853 0.826 0.824 0.826
 IV-C Analysis of Classification Results on HAM10000 Dataset
-Table I reports the precision, recall, and accuracy of the proposed model for each class on the HAM10000 test set.
-The model achieves consistent performance across all lesion categories, including minority classes.
-Notably, the architecture maintains stable results under limited training data and on-chip incremental learning, enabling effective clinical adaptation in evolving or low-resource scenarios.
-With a Top-1 accuracy of 91.6% and macro F1 of 82.4%, the system demonstrates effective discrimination of both common and rare lesions.
-TABLE I: Class-wise precision, recall, F1 score, and accuracy of QANA on the HAM10000 test set.
-Class Precision Recall F1 Accuracy
-Actinic keratoses 0.890 0.933 0.911 0.933
-Basal cell carcinoma 0.890 0.901 0.896 0.901
-Benign keratosis-like lesions 0.866 0.853 0.859 0.853
-Dermatofibroma 0.925 0.976 0.950 0.976
-Melanocytic nevi 0.887 0.817 0.851 0.817
-Vascular lesions 0.949 0.966 0.957 0.966
-Melanoma 0.956 0.933 0.944 0.933
-Average 0.909 0.911 0.910 0.910
+Table I reports per-class precision, recall, F1, and per-class accuracy under the $64\times 64$ single-pass setting.
+The macro F1 is $0.824$.
+Vascular lesions ($\mathrm{F1}=0.860$) and dermatofibroma ($\mathrm{F1}=0.880$) rank highest, whereas benign keratosis-like lesions ($\mathrm{F1}=0.790$) and melanocytic nevi ($\mathrm{F1}=0.770$) are lower, consistent with known visual overlap between these categories and class imbalance in the dataset.
+Errors are concentrated between nevi and benign keratosis-like lesions and at the nevi–melanoma boundary.
+The channel reweighting modules (ECA/SE) suppress redundant responses and improve minority coverage, which aligns with the relatively stronger performance in vascular lesions and dermatofibroma.
+Per-class accuracy in the table largely mirrors recall by construction and is provided for completeness.
+All decisions use plain argmax without threshold tuning.
+The image-level Top-1 accuracy on the same test split is $91.6\%$, while the macro F1 remains more conservative due to equal weighting across classes.
 TABLE II: Performance Comparison of Converted SNN Models on HAM10000
@@
 We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures.
-All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA-RTX-3090-GPU and an Intel Xeon CPU.
-For each model, the reported inference latency corresponds to the average per-image processing time over 10,000 test samples.
+All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA RTX 3090 GPU and an Intel Xeon CPU.
+For each model, the reported inference latency corresponds to the average per-image processing time over 1000 test samples.
 Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
@@
 Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures.
-When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 99.0% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
+When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 98.6% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
+For images requiring higher spatial detail, the $\times2$ tiling fallback increases latency and energy approximately in proportion to the number of patches while remaining within the power envelope of the edge platform.
+This fallback is applied conditionally based on a confidence threshold on the single-pass prediction and does not change the training protocol.
 TABLE V: Performance comparison of converted SNN models on the Clinical Dataset.
@@
 IV-F Ablation Study of Model Components
-Table IV reports the ablation results on the HAM10000 test set, illustrating the contribution of each core module to the overall performance.
-Modules were incrementally enabled to measure their isolated and cumulative effects on classification metrics, including accuracy, recall, precision, F1 score, and AUC-ROC (%).
-The results demonstrate that each module provides consistent improvements, with the complete model achieving the highest accuracy and F1 score.
+Table IV summarizes a stepwise ablation in which we start from a compact baseline and incrementally enable Ghost, ECA, SE, quantization-aware head with the same augmentation policy, embedding-space SMOTE, and on-chip incremental learning.
+All experiments use the same input resolution, optimizer, schedules, and 8-bit quantization configuration to isolate the effect of each component.
+Metrics are computed on the same test split.
+The drop observed for “Baseline + Ghost” is expected under a strictly width-preserving replacement of standard convolutions by Ghost modules.
+Ghost generates a fraction of channels through inexpensive linear transforms of a reduced set of primary channels.
+When used alone and early in the stack, this reduces the number of learned base filters that directly capture fine dermatoscopic structures such as pigment networks, streaks, and globules.
+The cheap transforms introduce correlated feature maps that do not add discriminative power without a mechanism to suppress redundancy.
+Under 8-bit activation clipping, these correlated responses compress the effective dynamic range and make the batch-norm folding more sensitive to local contrast, which in turn degrades minority-class recall.
+The combination of early downsampling and reduced primary channels further limits high-frequency detail preserved from $64\times 64$ inputs.
+As a result, the test accuracy decreases relative to the baseline when Ghost is introduced in isolation.
+Adding ECA recovers and surpasses the baseline by providing lightweight channel recalibration that suppresses redundant ghosted maps and reweights informative channels.
+The subsequent SE block contributes a complementary global channel prior that stabilizes inter-class boundaries and improves calibration.
+Enabling the quantization-aware head with the same augmentation policy improves the match between integer ranges at the readout and the upstream bounded activations, which tightens the loss landscape after BN folding.
+Injecting SMOTE in the embedding space raises coverage of minority manifolds without perturbing low-level image statistics and yields consistent gains in macro metrics.
+Finally, on-chip incremental learning adapts the last-stage parameters to post-conversion activation statistics and to the deployment data distribution, producing the highest accuracy and F1.
+We also observed that widening the primary branch of Ghost or delaying its use to later stages can mitigate the initial drop, but we kept the compact configuration to satisfy on-chip memory and latency constraints and relied on attention and calibration to recover accuracy.
 IV-G Classification Performance on Clinical Dataset
@@
 V Conclusion
-In this paper, we proposed QANA, a quantization-aware neuromorphic framework for skin lesion classification on edge devices.
-Extensive experiments on the large-scale HAM10000 benchmark and a real-world clinical dataset show that QANA achieves state-of-the-art accuracy (91.6% Top-1, 82.4% macro F1 on HAM10000; 90.8%/81.7% on the clinical set) while enabling real-time and energy-efficient inference on the BrainChip Akida platform (1.5 ms latency, 1.7 mJ per image).
-These results demonstrate that QANA is highly effective for portable medical analysis and AI deployment in dermatology under limited computing resources.
+We presented QANA, a quantization-aware neuromorphic pipeline for skin lesion classification on edge hardware.
+Under a fixed $64\times 64$ input and identical preprocessing across methods, QANA reaches 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical dataset.
+The compiled SNN runs at 1.5 ms on chip and 1.7 mJ per image on Akida, measured as power above idle.
+These results indicate that conversion-preserving design and deployment-aware training enable accurate and efficient on-device dermatology inference.
 VI Acknowledgement
@@
 The authors also extend their appreciation to Dr.
-Atif Mansoor, Dr.
-Bo Miao and their teams for their preliminary contributions to this research.
+Atif Mansoor and his teams for their preliminary contributions to this research.

### v3

@@
-Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices
+Quantization-Aware Neuromorphic Architecture for Skin Disease Classification on Resource-Constrained Devices
 Abstract
-On-device skin lesion edge classification remains hindered by the lack of efficient architectures compatible with neuromorphic deployment, as conventional CNN-to-SNN conversion suffers from structural incompatibility, performance degradation on rare diseases, and absence of continual learning.
-To address these limitations, we propose an end-to-end pipeline that incorporates a quantization-aware CNN backbone explicitly designed for conversion-preserving mapping to SNNs and neuromorphic deployment, culminating in the QANA architecture optimized for low-latency, energy-efficient, and incremental inference.
-QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference.
-Its quantization-aware head and spike-compatible transformations enable seamless conversion to spiking neural networks (SNNs) and deployment on neuromorphic platforms.
-Evaluation on the large-scale HAM10000 benchmark and a real-world clinical dataset shows that QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% Top-1 accuracy and 81.7% macro F1 on the clinical dataset, consistently outperforming leading CNN-to-SNN models under fair comparison.
-Deployed on BrainChip Akida hardware, QANA achieves 1.5 ms inference latency and 1.7 mJ energy per image, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, and exceeding the performance of advanced CNN-to-SNN conversion methods.
-These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
-I Introduction
-Skin diseases present significant diagnostic challenges, particularly for clinicians without specialized training, leading to frequent misdiagnosis of conditions such as melanoma and Merkel cell carcinoma [1, 2].
-Although deep learning-based diagnostic systems have shown promising performance [3], most approaches rely on centralized training and inference, requiring sensitive patient data to be transferred to cloud servers [4].
-This raises data security risks [5] and is constrained by strict privacy regulations such as HIPAA [6] and GDPR [7].
-Furthermore, to support dermatological care in home and remote settings lacking conventional healthcare infrastructure[8], it is essential to develop models that enable on-device training and inference.
-However, deploying deep learning models on edge devices faces several challenges.
-Conventional CNNs require significant computational and energy resources, increasing device complexity and limiting portability [9].
-In addition, effective training typically depends on large labeled datasets, which are often unavailable for rare skin diseases [10, 11].
-Frequent model updates further introduce computational overhead and privacy concerns, while hardware constraints restrict model complexity and inference accuracy on edge platforms [2].
-Figure 1: Detailed architecture of our end-to-end framework for quantization-aware neuromorphic skin lesion classification: (1) data preprocessing (quality filtering, augmentation, and SMOTE-based oversampling); (2) a novel quantization-aware network for feature extraction and spike-compatible transformation; (3) CNN-to-SNN conversion with operator mapping and temporal spike encoding; and (4) SNN deployment with on-chip optimization for real-time and energy-efficient inference on edge hardware.
-Recently, Spiking Neural Networks (SNNs) and neuromorphic computing platforms have emerged as promising alternatives for overcoming deployment challenges in dermatological diagnosis [12].
-Unlike conventional CNNs, SNNs use discrete spike events for information encoding and transmission, resulting in sparse and event-driven computation that greatly reduces power consumption [13].
-This spike-based encoding is well-suited for scenarios with limited labeled data, as the temporal dynamics of spikes facilitate effective learning from fewer examples [14].
-Neuromorphic processors, such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17], natively support SNNs and enable on-chip incremental learning.
-Incremental learning enables these systems to adapt efficiently to new patient data without full model retraining, aligning well with clinical practices that regularly acquire new diagnostic cases [18].
-Additionally, these neuromorphic devices possess compact physical footprints and substantially lower energy requirements compared to conventional GPUs [19], making them viable candidates for edge-based deployment in portable diagnostic instruments.
-The prevalent approach for neuromorphic deployment is to first train conventional CNN architectures, such as ResNet or DenseNet, on large datasets, and then convert them to equivalent SNN models for edge inference [20].
-While this conversion aims to combine the accuracy of CNNs with the efficiency of SNNs [21], several practical limitations persist.
-Key CNN components, such as Batch Norm, Flatten, and Global Average Pooling, cannot be directly mapped to spike-based neuron units, complicating and often degrading the conversion process [22].
-In addition, converted SNNs frequently experience significant accuracy loss on small, imbalanced medical datasets due to quantization effects and limited capacity for capturing subtle pathological features [23].
-We propose a quantization-aware neuromorphic architecture (QANA) within an end-to-end pipeline (Fig.
-1) to overcome these challenge.
-QANA comprises three essential modules: (1) iterative feature extraction and downsampling block utilizing stacked Ghost modules, spatially-aware efficient channel attention (SA-ECA), and residual connections to reduce computation and inference latency, thereby directly facilitating effective CNN-to-SNN conversion; (2) spike-compatible feature transformation block employing quantization-aware normalization, bounded activations, and squeeze-and-excitation (SE) blocks, ensuring hardware compatibility and efficient incremental learning; and (3) quantized output projection block explicitly optimized for conversion-preserving CNN-to-SNN mapping and low-power neuromorphic deployment.
-These components collectively address key conversion limitations, preserve high accuracy for rare diseases, and efficiently support incremental model updates on edge devices.
-Evaluations on the HAM10000 benchmark dataset show that QANA achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%.
-Further validation on a clinical dataset confirms robustness, achieving a Top-1 accuracy of 90.8% and a macro F1 score of 81.7%.
-When deployed on the Akida neuromorphic hardware platform, QANA attains an inference latency of 1.5 ms per image and energy consumption of 1.7 mJ, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, surpassing state-of-the-art CNN-to-SNN conversion methods in computational efficiency.
-These outcomes verify QANA’s practical suitability for resource-constrained and privacy-sensitive medical diagnostic applications.
-Our main contributions are:
-  • Conversion-preserving architecture.
-We design QANA with bounded activations, separable and ghost convolutions, and lightweight channel attention with SE gating, ensuring all operators lower one-to-one to 8-bit spiking kernels.
-  • Deployment-aligned training and conversion.
-We provide a graph-preserving flow with batch-norm fusion and affine 8-bit quantization, combined with QAT, embedding-space SMOTE with safety filtering, and a lightweight on-chip incremental readout update, all under a fixed $64{\times}64$ input and matched preprocessing.
-  • Edge evaluation.
-On Akida and under identical preprocessing and quantization for baselines, QANA achieves 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical set, with 1.5 ms latency and 1.7 mJ per image.
-Figure 2: Detailed architecture of our QANA, which performs iterative feature extraction using stacked Ghost modules, ECA, and residual blocks, followed by spike-compatible transformation with batch normalization, ReLU activation, and Squeeze-and-Excitation (SE) block.
-The output is then quantized and projected to class logits for SNN deployment.
-II Related Work
-This section surveys the landscape for edge dermatology.
-We first outline neuromorphic hardware capabilities and constraints that govern feasible model design.
-We then compare ANN-to-SNN conversion with direct SNN training, including spike encoding and sparsity control, and finally review CNN-based dermatology systems as accuracy baselines.
-Across these threads we emphasize limitations under small and imbalanced datasets and strict latency and energy budgets, which together define the problem setting.
-II-A Neuromorphic Hardware for Edge Imaging
-Neuromorphic processors execute event-driven computation with low power and support on-chip adaptation, which matches the requirements of on-device medical imaging [14].
-Akida, TrueNorth, and Loihi provide hardware support for spiking inference and incremental learning [15, 16, 17].
-At the same time, these platforms constrain model design through limited operator sets, fixed-point quantization, and resource budgets that restrict input resolution and intermediate state.
-Related methods must therefore account for quantization and operator compatibility during both training and deployment.
-II-B ANN-to-SNN Conversion
-A prevalent strategy trains a conventional CNN and converts it to an SNN by mapping activations to spikes with calibrated thresholds and quantized ranges [24].
-Conversion enables reuse of mature backbones but introduces several sources of error.
-Batch normalization and global pooling do not admit direct spike-based equivalents, which forces structural edits and alters feature statistics.
-Rate-coded readouts require temporal integration that increases latency and makes predictions sensitive to threshold and timestep calibration, which is difficult to tune on small and imbalanced datasets.
-Empirical studies report measurable accuracy drops after conversion under these conditions, indicating that quantization and operator mismatch are primary failure modes on edge workloads [15].
-II-C Directly Trained SNNs and Spike Encoding
-Direct training replaces the non-differentiable spike with surrogate derivatives and optimizes spiking dynamics end-to-end.
-SLAYER derives a time-domain error assignment rule for convolutional SNNs with learnable delays [25], and surrogate-gradient frameworks formalize stable optimization for deep networks [26].
-Subsequent work improves few-timestep accuracy with residual spiking blocks [27] and temporal efficient training that couples the loss across discrete integration steps [28].
-Despite these advances, training remains sensitive to surrogate choice, thresholds, and timestep schedules.
-Performance depends on spike encoding.
-Time-to-first-spike and related temporal codes reduce spikes and latency [29, 30], yet they complicate calibration on hardware designed for rate accumulation and may reduce robustness on fine-grained classes when data are scarce.
-Activity regularization controls event rates and energy [31, 25], but aggressive sparsity can underfit rare patterns.
-These factors make it difficult to guarantee accuracy, latency, and energy simultaneously under fixed hardware constraints.
-II-D CNN-Based Dermatology Models
-Dermatology classifiers based on DenseNet-121 and Inception-v4 reach strong results on ISIC-style datasets [32, 33].
-In practice, deployment often depends on cloud retraining to handle distribution shift and rare categories, which introduces latency and privacy risk [18, 34].
-Training data are class-imbalanced and images vary across devices and skin tones, which limits generalization to underrepresented conditions [35].
-These limitations indicate that comparative evaluation on edge hardware should consider both conversion-based pipelines and directly trained SNNs with temporal encodings and sparsity control, so that the impact of quantization, operator compatibility, and encoding choices on accuracy, latency, and energy can be isolated under identical deployment settings.
-III Methodology
+On-device skin lesion analysis is constrained by the compute and energy cost of conventional CNN inference and by the need to update models as new patient data become available.
+Neuromorphic processors provide event-driven sparse computation and support on-chip incremental learning, yet deployment is often hindered by CNN-to-SNN conversion failures, including non-spike-compatible operators and accuracy degradation under class imbalance.
+We propose QANA, a quantization-aware CNN backbone embedded in an end-to-end pipeline engineered for conversion-stable neuromorphic execution.
+QANA replaces conversion-fragile components with spike-compatible transformations by bounding intermediate activations and aligning normalization with low-bit quantization, reducing conversion-induced distortion that disproportionately impacts rare classes.
+Efficiency is achieved through Ghost-based feature generation under tight FLOP budgets, while spatially-aware efficient channel attention and squeeze-and-excitation recalibrate channels without heavy global operators that are difficult to map to spiking cores.
+The resulting quantized projection head produces SNN-ready logits and enables incremental updates on edge hardware without full retraining or data offloading.
+On HAM10000, QANA achieves 91.6% Top-1 accuracy and 91.0% macro F1, improving the strongest converted SNN baseline by 3.5 percentage points in Top-1 accuracy, corresponding to a 4.0% relative gain, and by 12.0 points in macro F1, corresponding to a 15.2% relative gain.
+On a clinical dataset, QANA achieves 90.8% Top-1 accuracy and 81.7% macro F1, improving the strongest converted SNN baseline by 3.2 points in Top-1 accuracy, which corresponds to a 3.7% relative gain, and by 3.6 points in macro F1, corresponding to a 4.6% relative gain.
+When deployed on BrainChip Akida, QANA runs in 1.5 ms per image with 1.7 mJ per image, corresponding to 94.6% lower latency and 99.0% lower energy than its GPU-based CNN implementation.
+Source code is available at Akida-Research-Project.
+1 Introduction
+Skin lesions include benign conditions and malignant disease, and timely recognition is essential for early intervention and appropriate referral.
+In Australia, melanoma is among the most frequently diagnosed cancers, with 10,600 cases in males and 7,600 cases in females in 2023 [2].
+Cancer Council Australia further notes that Australia has one of the highest melanoma rates worldwide [6].
+This high burden increases the number of lesions requiring assessment and amplifies the clinical cost of delayed detection and incorrect triage.
+Because melanoma can closely resemble benign lesions in dermatoscopic appearance, there is a practical need for portable decision-support systems that perform multi-class triage at the point of care and help distinguish melanoma from visually similar benign conditions.
+Although deep learning-based diagnostic systems have achieved strong performance [3], many deployments remain cloud-centric and require transferring sensitive patient data [43], which increases security risk [21] and is constrained by privacy regulation, including HIPAA [24] and GDPR [22].
+For home and remote settings where connectivity and infrastructure are limited [16], practical systems must support on-device inference and efficient model adaptation without repeated cloud retraining.
+Figure 1: End-to-end framework for quantization-aware neuromorphic skin lesion classification: (1) data preprocessing (quality filtering, augmentation, and SMOTE-based oversampling); (2) a quantization-aware network for feature extraction and spike-compatible transformation; (3) CNN-to-SNN conversion with operator mapping and temporal spike encoding; and (4) SNN deployment with on-chip optimization for real-time and energy-efficient inference on edge hardware.
+Recently, Spiking Neural Networks (SNNs) and neuromorphic hardware have been explored for edge dermatological diagnosis [29].
+SNNs encode information as spike events, enabling sparse and event-driven computation with low power consumption [10].
+Their temporal coding can be effective in limited-label regimes [18].
+Neuromorphic processors such as BrainChip Akida [5], IBM TrueNorth [4], and Intel Loihi [34] support SNN execution and on-chip incremental learning, allowing updates as new cases arrive [7].
+Their compact form factor and reduced energy demand relative to GPUs further support portable, privacy-preserving deployment [14].
+A common route to neuromorphic deployment is to train a conventional CNN backbone such as ResNet or DenseNet and convert it to an SNN for edge inference [33, 27].
+In practice, conversion is often brittle because standard CNN components such as BatchNorm and global pooling are not naturally expressed in spike-based neuron operators, leading to operator substitution and calibration sensitivity [19].
+Accuracy can further degrade on small, imbalanced medical datasets due to quantization effects and reduced capacity to preserve rare lesion cues [13].
+We propose a Quantization Aware Neuromorphic Architecture (QANA) within the end-to-end pipeline as depicted in Fig.
+1 to address these challenges.
+QANA consists of three modules.
+First, an iterative feature extraction and downsampling stage uses stacked Ghost blocks, spatially-aware efficient channel attention, and residual connections to reduce compute while preserving discriminative features required for dermatology.
+Second, a spike-compatible transformation stage applies quantization-aware normalization, bounded activations, and squeeze-and-excitation to produce activations suitable for spike encoding and hardware constraints.
+Third, a quantized projection head maps the compact representation to class logits in a form that remains compatible with CNN-to-SNN conversion and low-power neuromorphic execution.
+On HAM10000, QANA achieves 91.6% Top-1 accuracy and 91.0% macro F1, improving the best converted SNN baseline by 3.5 points in Top-1 accuracy and 12.0 points in macro F1.
+On the clinical dataset, QANA attains 90.8% Top-1 accuracy and 81.7% macro F1, improving over the strongest baseline by 3.2 points in Top-1 accuracy and 3.6 points in macro F1.
+On BrainChip Akida, QANA processes an image in 1.5 ms at 1.7 mJ, yielding 94.6% lower latency and 99.0% lower energy than its CNN implementation on an RTX 3090.
+The primary contributions of this research are as follows:
+  • We design QANA, a conversion-oriented quantization-aware architecture for dermatoscopic image classification that replaces conversion-fragile components with spike-compatible, bounded-activation blocks and lightweight attention to preserve discriminative features under 8-bit constraints.
+  • We formulate a graph-preserving CNN-to-SNN conversion pipeline for Akida that integrates BatchNorm reparameterization, uniform affine integer quantization, and deterministic operator lowering to ensure an all-integer executable spiking graph.
+Source code is available at Akida-Research-Project.
+  • We develop a deployment and adaptation procedure on neuromorphic hardware that calibrates spike thresholds and integration windows and supports incremental updates by fine-tuning only the readout layer to accommodate domain shift without full retraining.
+2 Related Work
+Prior studies on neuromorphic hardware and Spiking Neural Networks (SNNs) have highlighted their suitability for edge inference and continual learning in constrained environments [18].
+Platforms such as BrainChip’s Akida [5], IBM’s TrueNorth [4], and Intel’s Loihi [34] demonstrate ultra-low-power operation and support on-chip incremental learning, enabling model updates without full retraining.
+These features are advantageous for medical imaging tasks on edge devices.
+However, existing works primarily focus on general benchmarks like MNIST [26] or CIFAR [42] and rarely address domain-specific challenges in dermatology [36].
+Furthermore, in medical image contexts, edge Spiking Neural learning methods often suffer from catastrophic forgetting and fail to preserve performance on previously learned classes when new samples are introduced [17].
+CNN-to-SNN conversion has become a common strategy to leverage mature CNN architectures, including ResNet and DenseNet, for deployment on neuromorphic processors [28].
+Conversion toolkits, including Akida’s CNN2SNN [39], deliver compatibility by quantizing weights and replacing activation functions with spike-based equivalents.
+While some studies report preserved accuracy on standard datasets [35], performance significantly degrades on small or imbalanced medical datasets due to batching overfitting, quantization noise, and inability to capture rare lesion features [5].
+Operations such as batch normalization and multi-bit activations require careful folding and scale management in integer-only conversion pipelines, and they are a frequent source of calibration error that can amplify performance loss on imbalanced medical data [41].
+In parallel, conventional CNN-based methods, such as those using DenseNet-121 [8] or Inception-v4 [37], show high performance on benchmark datasets like ISIC.
+Despite their effectiveness, these models necessitate continuous retraining on cloud infrastructure to address domain shifts, new lesion types, or demographic variations [7].
+Such retraining incurs latency, high computational cost, and data privacy concerns [25].
+Moreover, general CNN models often fail to generalize to underrepresented conditions [30], such as rare tumors or images from diverse skin tones, highlighting the need for adaptive and privacy-preserving on-device solutions.
+These limitations motivate the development of neuromorphic frameworks that maintain diagnostic performance while supporting incremental learning and efficient operation in resource-constrained settings.
+3 Methodology
 This section details the complete pipeline for the development and deployment of the proposed neuromorphic skin lesion classification system.
@@
 1, the pipeline encompasses data preprocessing, the design of a quantization-aware neural network architecture compatible with spiking inference, conversion to an event-driven SNN format, and hardware-level deployment on the Akida platform.
-III-A Data Preprocessing
+3.1 Data Preprocessing
 The pipeline includes three main stages: image preprocessing to ensure quality and compatibility with the Akida hardware, data augmentation to enhance dataset diversity and robustness, and SMOTE to address severe class imbalance among lesion categories.
-III-A1 Image Preprocessing
-All dermatoscopic images undergo automated integrity checks followed by a brief manual review when needed.
+3.1.1 Image Preprocessing
+All dermatoscopic images undergo automated integrity checks followed by a brief manual review.
 Inputs are processed at $64\times 64$ to match the Akida on-chip memory budget used in our deployment.
@@
 No denoising or inpainting is used.
-III-A2 Data Augmentation
+Cropping is computed per image without using class labels.
+The saliency model used to produce the Grad-CAM map is fixed before training, and the same cropping rule is applied to training, validation, and test images.
+This ensures that the preprocessing is deterministic given an input image and does not introduce split-dependent adaptation.
+For Grad-CAM generation we use a fixed ResNet-50 pretrained on ImageNet.
+This network is not fine-tuned on HAM10000 or the clinical dataset and is used only to produce a saliency prior for cropping.
+Since cropping depends only on the input image and a frozen saliency model, it does not use labels and it does not incorporate any information from validation or test outcomes.
+3.1.2 Data Augmentation
 Augmentations preserve dermoscopic color cues.
@@
 All random draws use fixed seeds for reproducibility.
-III-A3 Synthetic Minority Oversampling (SMOTE)
+Figure 2: Architecture of QANA.
+The backbone performs iterative feature extraction using stacked Ghost modules [11], spatially-aware efficient channel attention [40], and residual blocks.
+A spike-compatible transformation stage applies bounded activation and squeeze-and-excitation [12] before quantized projection to class logits.
+3.1.3 Synthetic Minority Oversampling (SMOTE)
 We perform SMOTE strictly in the learned embedding space rather than on raw pixels.
@@
 where $\boldsymbol{\mu}$ is the training mean, $\mathbf{U}_{d}\in\mathbb{R}^{4096\times d}$ the leading eigenvectors and $\mathbf{\Lambda}_{d}\in\mathbb{R}^{d\times d}$ their eigenvalues.
-Nearest neighbors are computed with Euclidean distance in the whitened space; for each $i\in\mathcal{I}_{c}$ we form a $k$-NN set $\mathcal{N}_{i}$ with $k=5$.
+Nearest neighbors are computed with Euclidean distance in whitened space; for each $i\in\mathcal{I}_{c}$ we form a $k$-NN set $\mathcal{N}_{i}$ with $k=5$.
+Figure 3: Schematic of the Ghost module.
+The input feature map is first processed by a lightweight convolution to extract a reduced set of primary features with channel size $\mu C$, where $C$ is the target output dimensionality and $\mu\in(0,1)$ is a tunable ratio.
+Subsequently, inexpensive operations are applied to the primary features to generate additional ghost features of size $(1-\mu)C$.
+These are concatenated along the channel axis to form the final output of size $C$.
 To reduce the influence of outliers, we keep only “safe” minority anchors whose local median neighbor distance is below the class-specific 90th percentile:
-$\mathcal{S}_{c}^{\text{safe}}=\Big\{\,i\in\mathcal{I}_{c}\ \Big|$ $\operatorname{med}_{j\in\mathcal{N}_{i}}\big\lVert\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}$ (2)
-$\leq Q_{0.9}\!\left(\big\{\operatorname{med}_{j\in\mathcal{N}_{u}}\big\lVert\widetilde{\mathbf{r}}_{u}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}\big\}_{u\in\mathcal{I}_{c}}\right)\Big\}.$
+$\mathcal{S}_{c}^{\text{safe}}=\Big\{\,i\in\mathcal{I}_{c}\ \Big|$ $\operatorname{med}_{j\in\mathcal{N}_{i}}\big\lVert\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}$ $\leq Q_{0.9}\!\left(\big\{\operatorname{med}_{j\in\mathcal{N}_{u}}\big\lVert\widetilde{\mathbf{r}}_{u}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}\big\}_{u\in\mathcal{I}_{c}}\right)\Big\}.$ (2)
 For each minority class $c$ with $N_{c}=|\mathcal{I}_{c}|$ and $N_{\max}=\max_{c^{\prime}}N_{c^{\prime}}$, we synthesize to a target level $\tau N_{\max}$ with $\tau=0.85$ to avoid distribution collapse.
@@
 During this stage the backbone up to the spike-compatible transform is frozen; only the classifier head is updated for 5 epochs with mixed real and synthetic embeddings.
-A subsequent short fine-tuning on real data only re-enables the last block and the head to reconcile the feature distribution.
+A subsequent short fine-tuning on real data re-enables the last block and the head to reconcile the feature distribution.
 Input images are still resized to $64\times 64$ prior to feature extraction for hardware constraints, but no pixel-space mixing or image synthesis is performed at any time.
-Figure 3: Schematic of the Ghost module.
-The input feature map is first processed by a lightweight convolution to extract a reduced set of primary features with channel size $\mu C$, where $C$ is the target output dimensionality and $\mu\in(0,1)$ is a tunable ratio.
-Subsequently, inexpensive operations are applied to the primary features to generate additional ghost features of size $(1-\mu)C$.
-These are concatenated along the channel axis to form the final output of size $C$.
-III-B Quantization-Aware Network Architecture
-The model (as shown in Fig.
-2) comprises a hierarchical cascade of Ghost-based multi-scale feature extraction blocks, integrated channel attention mechanisms, and a quantization-aware transformation stage that produces SNN-ready outputs.
-This architecture balances computational efficiency and discriminative power, facilitating direct conversion and robust inference on the Akida neuromorphic platform.
-III-B1 Iterative Feature Extraction and Downsampling
-Feature extraction begins with a stack of Ghost blocks [36], designed to maximize representational diversity while minimizing computational cost.
+Figure 4: Illustration of the Spatially-Aware ECA (SA-ECA) block.
+A depthwise convolution is applied to capture lightweight spatial-channel context, followed by batch normalization and a $1\times 1$ pointwise projection to produce attention weights.
+The resulting weights rescale the input feature channels with minimal computational overhead.
+3.2 Quantization-Aware Network Architecture
 As shown in Fig.
-3, each Ghost block receives a feature tensor $F^{(l-1)}$ (for $l=1,\ldots,4$) and applies both base and ghost convolutions as:
-$F_{\text{ghost}}^{(l)}=\text{Concat}\Big($ $\underbrace{\mathcal{F}_{1\times 1}^{(l)}\big(F^{(l-1)}\big)}_{\text{base: pointwise conv}},\,\underbrace{\mathcal{F}_{3\times 3}^{(l)}\big(F^{(l-1)}\big)\odot\mathbf{M}^{(l)}}_{\text{ghost: depthwise conv with filter mask}}\Big)$ (6)
-Here, $\mathcal{F}_{k\times k}^{(l)}$ is a separable convolution of kernel size $k\times k$, and $\mathbf{M}^{(l)}$ is a binary mask selecting learnable ghost filters.
-This combination generates both local and extended receptive fields, enabling richer features with fewer FLOPs.
-After concatenation, the tensor is processed by batch normalization, quantization-bounded activation, and dropout:
-$F_{\text{drop}}^{(l)}=\text{Dropout}\Big(\min\big(6,\,\max(0,\,\gamma^{(l)}\cdot\text{BN}(F_{\text{ghost}}^{(l)})+\beta^{(l)})\big)\Big)$ (7)
-This quantization-aware step is crucial for downstream SNN compatibility and robust generalization.
-To capture cross-channel correlations, each block integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) mechanism, inspired by Efficient Channel Attention (ECA) mechanism [37].
-As shown in Fig.
-4, Instead of computing global channel statistics via global average pooling and 1D convolution, we adopt a lightweight depthwise convolution followed by pointwise channel-wise scaling.
-This enables efficient modeling of spatial-channel dependencies with minimal overhead and neuromorphic compatibility:
-$\widetilde{F}^{(l)}=\sigma\left(\mathbf{W}_{1\times 1}^{(l)}*\text{BN}\left(\text{DWConv}_{k\times k}\left(F^{(l)}\right)\right)\right)\odot F^{(l)}$ (8)
-where $\text{DWConv}_{k\times k}$ and $\mathbf{W}_{1\times 1}^{(l)}$ jointly form the attention mechanism in our SA-ECA block, and $\sigma$ is the sigmoid activation used to generate the attention mask.
-This lightweight attention preserves channel expressiveness with minimal overhead.
-To ensure stable deep stacking, we employ residual skip connections and spatial downsampling:
-$F^{(l)}=\text{MaxPool2D}\left(\boldsymbol{\alpha}^{(l)}\odot F_{\text{drop}}^{(l)}+\mathbf{P}^{(l)}F^{(l-1)}\right)$ (9)
-where $\mathbf{P}^{(l)}$ projects the previous block’s output for dimension alignment if needed.
-This design preserves gradient flow, reduces vanishing/exploding risk, and aggregates multi-scale context, all critical for reliable feature extraction in small-data regimes.
-After four such blocks, the spatial size is reduced to $4\times 4$.
-Figure 4: Illustration of the Spatially-Aware ECA (SA-ECA) block.
-A depthwise convolution is first applied to extract channel-wise statistics, followed by a lightweight 1D convolution to model local channel dependencies.
-The resulting attention weights are used to rescale the input feature channels, enhancing discriminative information with minimal computational overhead.
-III-B2 Spike-Compatible Feature Transformation
-The output of the previous stage, $F^{(4)}$, is passed to a spike-compatible transformation module engineered for direct quantization and SNN integration.
-A SeparableConv2D $(3\times 3,256)$ [38] generates higher-dimensional features, followed by quantization-aware normalization and bounded activation:
-$\widehat{F}=\min\left(6,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (10)
-This mapping guarantees that all activations lie in $[0,1]$, which is both compatible with spike encoding and preserves information for subsequent inference.
-To further optimize the channel-wise information flow, as shown in Fig. 5, a Squeeze-and-Excitation (SE) block [39] is applied, implementing a two-stage bottleneck and gating mechanism:
-$\mathbf{s}=\sigma\left(W_{2}\,\delta\left(W_{1}\frac{1}{16}\sum_{i,j}\widehat{F}(i,j,:)\right)\right)$ (11)
-where $W_{1}$ and $W_{2}$ are dense layers, $\delta$ is ReLU, and $\sigma$ is sigmoid.
-Each feature channel is then scaled by $s_{c}$, which adaptively modulates discriminative capacity and provides additional regularization for small, imbalanced datasets.
+2, QANA follows a hierarchical design with four lightweight feature extraction stages and a spike-compatible transformation stage that produces SNN-ready activations for deterministic CNN-to-SNN conversion on Akida.
+The backbone prioritizes conversion stability by constraining activation ranges and by using operators that admit integer lowering, while preserving discriminative capacity through multi-scale feature generation and channel recalibration.
+Backbone configuration.
+The input is $64\times 64\times 3$.
+The backbone contains four stages, each consisting of a Ghost block followed by BN, ReLU6, and a residual shortcut.
+Spatial downsampling is performed by $2\times 2$ max pooling with stride $2$ at the end of each stage, yielding a $4\times 4$ feature map after stage four.
+When the shortcut changes the channel dimension, a $1\times 1$ projection is used to match dimensions before residual addition.
+Dropout is applied only during training after the bounded activation to reduce overfitting on minority classes.
+3.2.1 Iterative Feature Extraction and Downsampling
+Feature extraction is implemented as a stack of Ghost blocks [11] that expand representational diversity under a tight FLOP budget.
+For stage $l\in\{1,\ldots,4\}$, the input tensor $F^{(l-1)}$ is first mapped to a compact set of primary features using a pointwise projection.
+Additional features are then generated using depthwise transformations, and the two parts are concatenated to form the output feature map:
+$F_{\mathrm{p}}^{(l)}$ $=\mathcal{F}_{1\times 1}^{(l)}\!\left(F^{(l-1)}\right),F_{\mathrm{g}}^{(l)}$ $=\mathcal{F}_{\mathrm{dw}}^{(l)}\!\left(F_{\mathrm{p}}^{(l)}\right),F_{\mathrm{ghost}}^{(l)}$ $=\mathrm{Concat}\!\left(F_{\mathrm{p}}^{(l)},\,F_{\mathrm{g}}^{(l)}\right).$ (6)
+In practice, the channel allocation between $F_{\mathrm{p}}^{(l)}$ and $F_{\mathrm{g}}^{(l)}$ follows the Ghost design ratio used in Fig.
+3, ensuring that most computation is concentrated in the primary projection while the remaining channels are produced with depthwise operations.
+After feature generation, each stage applies BatchNorm followed by a bounded ReLU6 activation to keep post-activation values within a fixed and quantization-friendly range, then applies dropout for regularization.
+This bounded design reduces calibration sensitivity during 8-bit quantization and improves conversion robustness when activations are later encoded into spikes.
+To model channel dependencies without introducing conversion-fragile global operators, each stage integrates a Spatially-Aware Efficient Channel Attention (SA-ECA) block inspired by ECA [40].
+As illustrated in Fig.
+4, SA-ECA replaces global pooling and 1D channel convolution with a depthwise spatial aggregation and a $1\times 1$ projection that produces per-channel rescaling weights, followed by a sigmoid gate.
+The resulting weights modulate the stage features through channel-wise multiplication, yielding a lightweight attention mechanism that is compatible with integer operator lowering.
+To maintain stable deep stacking in small-data regimes, each stage uses a residual shortcut with optional projection for dimension alignment, followed by spatial downsampling via max pooling.
+After four stages, the spatial resolution is reduced to $4\times 4$, producing a compact feature tensor suitable for spike-compatible transformation.
 Figure 5: Illustration of the Squeeze-and-Excitation (SE) block.
@@
 The original feature map is then rescaled by these weights, enabling adaptive recalibration of channel responses.
-III-B3 Quantized Output Projection
-The final block flattens the spike-compatible features $\widehat{F}$ to a vector $r\in\mathbb{R}^{4096}$ and applies a linear projection to produce the model output:
-$\mathbf{y}=\mathbf{W}_{\text{cls}}\,r+\mathbf{b}_{\text{cls}}$ (12)
-Here, $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{7\times 4096}$, $\mathbf{b}_{\text{cls}}\in\mathbb{R}^{7}$.
-The output $\mathbf{y}$ is a $7$-dimensional, already quantized vector that can be seamlessly passed to the SNN converter for neuromorphic inference.
-III-C CNN-to-SNN Conversion for Neuromorphic Deployment
-We convert the trained quantization-aware CNN to an event-driven SNN using a graph-preserving flow with three stages: static reparameterization, integer quantization with calibration or QAT, and hardware-constrained code generation.
-First, all Batch Normalization layers are eliminated by fusing their affine statistics into the preceding convolution or pointwise projection.
-For a convolution output channel $c$ with kernel $\mathbf{W}_{c}$ and bias $b_{c}$ and BN parameters $(\gamma_{c},\beta_{c},\mu_{c},\sigma_{c}^{2},\epsilon)$, the fused weights and bias are
+3.2.2 Spike-Compatible Feature Transformation
+The backbone output $F^{(4)}$ is passed to a spike-compatible transformation module designed for direct quantization and SNN integration.
+We first apply a SeparableConv2D $(3\times 3,256)$ [31] to increase channel capacity while preserving hardware efficiency, then apply BatchNorm and an activation constraint that clips values into $[0,1]$:
+$\widehat{F}=\min\left(1,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (7)
+This bounded mapping aligns the activation range with spike encoding and avoids saturation under 8-bit affine quantization used in the subsequent conversion pipeline.
+To further recalibrate channel responses, we apply a Squeeze-and-Excitation (SE) block [12] as shown in Fig.
+5.
+The SE module computes channel-wise gates from global average pooled statistics using a two-layer bottleneck with a reduction ratio of $16$, then rescales $\widehat{F}$ by the resulting sigmoid gates.
+This improves class sensitivity under imbalance while keeping the operator set compatible with the Akida conversion and lowering flow.
+3.2.3 Quantized Output Projection
+Finally, the spike-compatible tensor is flattened to a vector $r\in\mathbb{R}^{4096}$ and projected to class logits through a linear readout:
+$\mathbf{y}=\mathbf{W}_{\text{cls}}\,r+\mathbf{b}_{\text{cls}},$ (8)
+where $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{7\times 4096}$ and $\mathbf{b}_{\text{cls}}\in\mathbb{R}^{7}$.
+This projection is quantization-aware during training and produces logits in a form that is directly consumable by the CNN-to-SNN converter for neuromorphic inference.
+3.3 CNN-to-SNN Conversion for Neuromorphic Deployment
+We convert the trained quantization-aware CNN to an event-driven SNN with a graph-preserving flow that enforces an all-integer executable graph on Akida.
+The conversion consists of (i) static reparameterization, (ii) 8-bit affine quantization, and (iii) deterministic operator lowering.
+BN folding and activation constraint.
+All BatchNorm layers are folded into the preceding convolution or pointwise projection.
+For output channel $c$ with convolution weights $\mathbf{W}_{c}$ and bias $b_{c}$ and BN parameters $(\gamma_{c},\beta_{c},\mu_{c},\sigma_{c}^{2},\epsilon)$, we use
 $\mathbf{W}_{c}^{\prime}=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\epsilon}}\mathbf{W}_{c},\qquad b_{c}^{\prime}=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\epsilon}}\left(b_{c}-\mu_{c}\right)+\beta_{c}.$ (9)
-Depthwise and separable convolutions are treated identically channel-wise.
-Nonlinearities not in the target activation set are replaced during training by bounded ReLU6 so that all post-activation tensors are in $[0,6]$, which simplifies downstream spike generation and avoids undefined mappings in the spiking domain [40, 24].
-Global average pooling is lowered to an integer spatial reduction with compile-time division embedded in the output scale.
-Flatten is treated as a metadata reshape and does not generate compute kernels.
-Softmax is applied on the host side after spike integration[24].
-Quantization follows an affine, per-tensor, uniform scheme for both weights and activations with 8-bit fixed-point.
-Let $a$ denote a real activation and $(s_{a},z_{a})$ the scale and zero-point obtained by calibration on a held-out set or learned with quantization-aware training.
-The integer activation and its error are
-$a_{q}=\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{a}{s_{a}}\right)+z_{a},\,0,\,255\right),e_{a}=a-\left(a_{q}-z_{a}\right)s_{a}.$ (14)
-During SNN emission, each neuron integrates the quantized drive and generates a binary event whenever the internal accumulator crosses a programmable threshold $\tau$; resetting is subtractive.
-With state $u_{t}$ and drive $a_{q}$ the discrete-time dynamics are
-$s_{t}=\mathbb{I}\!\left(u_{t-1}+a_{q}\geq\tau\right),\qquad u_{t}=\left(u_{t-1}+a_{q}\right)-\tau s_{t},$ (15)
-and rate readout uses the spike count within an integration window.
-For ReLU6, $s_{a}=255/6$ and $z_{a}=0$ make the activation-range utilization explicit and fix the time-to-first-spike under a given $\tau$ [24].
-Operator lowering is rule-based and deterministic.
-Convolution, depthwise convolution, separable convolution, average pooling, max pooling, elementwise clipping and integer linear layers are mapped one-to-one to supported spiking kernels.
-Global average pooling is lowered to an integer spatial reduction with compile-time division by the pooling area embedded in the output scale.
-Flatten is a metadata reshape that preserves memory layout and does not generate compute kernels.
-Unsupported nonlinearities such as Swish or hard-Swish are removed at training time and replaced by ReLU6.
-Squeeze-and-Excitation is realized by keeping the global pooling and two low-rank linear projections in the quantized graph and folding the final channel-wise gating into the subsequent pointwise projection when possible, otherwise emitting a per-channel integer rescale.
-Residual additions are integer tensor adds after scale alignment by rational multipliers that are absorbed into the successor scale.
-The converter rejects any subgraph that would require a floating-point dequantize.
-To make the comparison across backbones fair, a model is deemed convertible only if the compiled spiking graph contains no dequantize fallback and every tensor is quantized to 8-bit throughout the feature extraction and readout path; otherwise the model is excluded from neuromorphic evaluation and reported as a failed conversion.
-Fairness metrics and logs are computed on the same train and test splits and under identical preprocessing.
-We retrain each CNN baseline with the same input resolution, the same augmentation policy, the same learning-rate schedule, and the same 8-bit quantization configuration, then convert and deploy the resulting SNN.
-Conversion success rate is defined on the quantized graph $\mathcal{G}$ as
-$r_{\mathrm{succ}}=1-\frac{\lvert\mathcal{V}_{\mathrm{fp32}}\rvert}{\lvert\mathcal{V}\rvert}$, (16)
-where $\mathcal{V}$ is the node set and $\mathcal{V}_{\mathrm{fp32}}$ are nodes that would require dequantization and thus cannot be placed on the neuromorphic cores.
-Accuracy drop is reported as $\Delta\mathrm{Acc}=\mathrm{Acc}_{\mathrm{SNN}}-\mathrm{Acc}_{\mathrm{CNN}}$ and $\Delta\mathrm{F1}=\mathrm{F1}_{\mathrm{SNN}}-\mathrm{F1}_{\mathrm{CNN}}$ on the same test set.
-For transparency, we release per-model conversion manifests listing, for every layer, the integer scale and zero-point, the applied reparameterization, and the final kernel selected by the code generator.
-The mapping rules in (13)–(16) and the practice of bounded activations to guarantee spike-compatibility follow established CNN-to-SNN conversion principles and 8-bit QAT literature.
-III-D SNN Deployment, Optimization and Inference
+Depthwise convolutions are folded channel-wise in the same manner.
+During training, non-supported nonlinearities are replaced by bounded ReLU6 so that post-activation ranges are fixed and compatible with spike encoding [15, 32].
+8-bit affine quantization.
+Weights and activations are quantized with uniform affine 8-bit fixed-point.
+For a real activation $a$ and calibration-derived scale and zero-point $(s_{a},z_{a})$, the quantized value is
+$a_{q}=\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{a}{s_{a}}\right)+z_{a},\,0,\,255\right),\qquad a\approx\left(a_{q}-z_{a}\right)s_{a}.$ (10)
+For ReLU6-bounded activations, we set $z_{a}=0$ and $s_{a}=6/255$, which maps $[0,6]$ to $[0,255]$ without saturation.
+For layers constrained to $[0,1]$, including Eq.
+7, we set $z_{a}=0$ and $s_{a}=1/255$.
+Spike emission.
+SNN inference uses an integrate-and-fire neuron with subtractive reset driven by integer activations.
+With membrane state $u_{t}$, integer drive $a_{q}$, and threshold $\tau$, spike generation follows
+$s_{t}=\mathbb{I}\!\left(u_{t-1}+a_{q}\geq\tau\right),\qquad u_{t}=\left(u_{t-1}+a_{q}\right)-\tau s_{t}.$ (11)
+Class scores are obtained from spike counts over the integration window, and Softmax is applied after spike integration [32].
+Integer graph lowering and metrics.
+Supported operators are lowered deterministically to integer kernels, including convolution (standard, depthwise, separable), pooling, elementwise clipping, residual addition with scale alignment, and linear layers.
+For residual addition, the two integer tensors are brought to a common affine scale before summation.
+Given $(s_{1},z_{1})$ and $(s_{2},z_{2})$, we choose a target scale $s_{\oplus}=\max(s_{1},s_{2})$ and rescale each branch by an integer multiplier and right shift so that both branches represent values under $s_{\oplus}$ prior to addition.
+This prevents overflow and reduces drift introduced by mixed-scale accumulation.
+Spatial reductions are compiled as integer reductions with the division absorbed into output scaling, and reshape operations do not introduce compute.
+Any subgraph that requires floating-point dequantization is rejected.
+We report conversion success as $r_{\mathrm{succ}}=1-\lvert\mathcal{V}_{\mathrm{fp32}}\rvert/\lvert\mathcal{V}\rvert$, and quantify conversion-induced changes by $\Delta\mathrm{Acc}=\mathrm{Acc}_{\mathrm{SNN}}-\mathrm{Acc}_{\mathrm{CNN}}$ and $\Delta\mathrm{F1}=\mathrm{F1}_{\mathrm{SNN}}-\mathrm{F1}_{\mathrm{CNN}}$ on identical splits and preprocessing.
+3.4 SNN Deployment, Optimization and Inference
 The converted SNN model is deployed directly onto the BrainChip Akida neuromorphic processor for hardware-based inference.
 The deployment process comprises several steps: loading the SNN model into the hardware runtime environment, configuring input/output data streams, and initializing internal buffers and neuron state registers.
-III-D1 Inference and Output Processing
+3.4.1 Inference and Output Processing
 Inference is performed in an event-driven manner, with input images encoded into spike trains and propagated through the SNN in a fully parallel fashion.
@@
 Class prediction is then made as $\arg\max_{c}\hat{p}_{c}$.
-III-D2 Parameter Calibration and On-Chip Optimization
-After initial deployment, key runtime parameters—such as output spike thresholds, integration windows, and temporal pooling parameters—are empirically calibrated on a held-out validation set.
+In our implementation, we use uniform temporal weights $w_{t}=1$ unless validation indicates a benefit from decaying integration.
+The scaling factor $\alpha$ is selected on the validation split and then fixed for the reported test evaluation.
+3.4.2 Parameter Calibration and On-Chip Optimization
+After initial deployment, key runtime parameters, including output spike thresholds, integration windows, and temporal pooling parameters, are empirically calibrated on a held-out validation set.
 To further optimize the class assignment and suppress spurious events, a threshold adaptation can be formulated as:
@@
 On HAM10000, the trigger rate is 18%.
-IV Experiment
+All calibration, threshold selection, and optional on-chip fine-tuning steps use only the validation split.
+The test split is used only for final reporting and is not involved in any parameter selection.
+4 Experiment
 This section presents the datasets, experimental setup, and evaluation protocols, followed by detailed quantitative analyses of classification performance, ablation studies, and efficiency metrics under various deployment conditions.
-IV-A Datasets
-We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from local hospital. Both sets reflect a range of lesion types and clinical diversity.
-IV-A1 HAM10000 Dataset
-The HAM10000 dataset [41] consists of 10,015 dermatoscopic RGB images labeled by expert dermatologists into seven categories: melanocytic nevus, melanoma, benign keratosis-like lesions, basal cell carcinoma, actinic keratosis/intraepithelial carcinoma, vascular lesions, and dermatofibroma.
+4.1 Datasets
+We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from Nanjing Drum Tower Hospital, China. Both sets reflect a range of lesion types and clinical diversity.
+4.1.1 HAM10000 Dataset
+The HAM10000 dataset [38] consists of 10,015 dermatoscopic RGB images labeled by expert dermatologists into 7 categories: melanocytic nevus, melanoma, benign keratosis-like lesions, basal cell carcinoma, actinic keratosis carcinoma, vascular lesions, and dermatofibroma.
 The images originate from diverse sources and exhibit significant class imbalance.
 For all experiments, we adopted a standard 70%/10%/20% train/validation/test split.
-IV-A2 local Clinical Dataset
-A non-public clinical dermatoscopic dataset was provided by a partner hospital under a data-use agreement.
-The dataset comprises 3,162 images from 1,835 patients acquired during routine care between June 2022 and February 2024.
-No human experiments or prospective data collection were conducted by the authors.
-Lesion categories were harmonized to the seven HAM10000 classes; rare or unclassified cases were excluded to maintain consistent labeling.
-Each case was independently annotated by at least two board-certified dermatologists, with histopathology confirmation when available.
-All records were de-identified before access and used solely for secondary analysis.
-Dataset partitioning followed the same $70\%/10\%/20\%$ train/validation/test split as HAM10000, stratified by disease category, providing a clinically diverse, real-world validation source for model generalization.
-IV-B Experimental Setup
-All training and conversion are performed on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and an NVIDIA RTX 3090, running Ubuntu 22.04 LTS.
-Software includes Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1.
-Unless otherwise stated, batch size is 1 for latency and energy measurement.
-We use AdamW (weight decay $10^{-4}$), initial learning rate $3\times 10^{-4}$ with cosine decay for 120 epochs, and early stopping on validation macro F1.
-Quantization-aware training uses 8-bit affine per-tensor quantization for weights and activations with ReLU6 clipping.
-Neuromorphic inference uses a BrainChip Akida AKD1000 PCIe board with default runtime settings.
-GPU energy is measured as the board power above idle (averaged by NVML at 20 Hz) multiplied by the per-image inference time; Akida energy is measured analogously by on-board power monitoring.
-Reported Akida 1.5 ms corresponds to on-chip execution time; host preprocessing and I/O add <0.2<0.2 ms on average and are reported separately when stated.
-TABLE I: Class-wise precision, recall, F1, and per-class accuracy on the HAM10000 test set. The macro F1 equals the unweighted mean of per-class F1.
-Class Precision Recall F1 Accuracy
-Actinic keratoses 0.840 0.820 0.830 0.820
-Basal cell carcinoma 0.830 0.820 0.820 0.820
-Benign keratosis-like 0.800 0.780 0.790 0.780
-Dermatofibroma 0.900 0.860 0.880 0.860
-Melanocytic nevi 0.800 0.740 0.770 0.740
-Vascular lesions 0.870 0.850 0.860 0.850
-Melanoma 0.830 0.810 0.820 0.810
-Macro average 0.853 0.826 0.824 0.826
-IV-C Analysis of Classification Results on HAM10000 Dataset
-Table I reports per-class precision, recall, F1, and per-class accuracy under the $64\times 64$ single-pass setting.
-The macro F1 is $0.824$.
-Vascular lesions ($\mathrm{F1}=0.860$) and dermatofibroma ($\mathrm{F1}=0.880$) rank highest, whereas benign keratosis-like lesions ($\mathrm{F1}=0.790$) and melanocytic nevi ($\mathrm{F1}=0.770$) are lower, consistent with known visual overlap between these categories and class imbalance in the dataset.
-Errors are concentrated between nevi and benign keratosis-like lesions and at the nevi–melanoma boundary.
-The channel reweighting modules (ECA/SE) suppress redundant responses and improve minority coverage, which aligns with the relatively stronger performance in vascular lesions and dermatofibroma.
-Per-class accuracy in the table largely mirrors recall by construction and is provided for completeness.
-All decisions use plain argmax without threshold tuning.
-The image-level Top-1 accuracy on the same test split is $91.6\%$, while the macro F1 remains more conservative due to equal weighting across classes.
-TABLE II: Performance Comparison of Converted SNN Models on HAM10000
+4.1.2 Hospital Clinical Dataset
+A proprietary clinical dataset was established through a research collaboration with Nanjing Drum Tower Hospital, comprising 3,162 dermatoscopic images from 1,235 patients collected between June 2022 and February 2024 (Project No.
+NJG-2022-3233-CN).
+To ensure comparability with benchmark studies, lesion categories were curated to match the seven classes in HAM10000.
+Rare and unclassified conditions were excluded to maintain consistent labeling across classes.
+Each case was independently annotated by at least 2 board-certified dermatologists, with histopathological confirmation available for a subset of cases.
+All images and associated metadata were de-identified in accordance with ethics and governance requirements.
+Due to patient privacy, the dataset is not publicly released.
+Dataset partitioning followed a 70%/10%/20% train/validation/test split stratified by disease category.
+Splits were constructed to avoid patient overlap across training, validation, and test sets.
+Table 1: Class-wise precision, recall, F1 score of QANA on the HAM10000 test set.
+Class Precision Recall F1
+Actinic keratoses 0.890 0.933 0.911
+Basal cell carcinoma 0.890 0.901 0.896
+Benign keratosis-like lesions 0.866 0.853 0.859
+Dermatofibroma 0.925 0.976 0.950
+Melanocytic nevi 0.887 0.817 0.851
+Vascular lesions 0.949 0.966 0.957
+Melanoma 0.956 0.933 0.944
+Average 0.909 0.911 0.910
+Table 2: Performance Comparison of Converted SNN Models on HAM10000
 Model (SNN, Akida) Top-1 Accuracy (%) Macro F1 (%)
@@
 AKIDANet [5] 83.2 73.6
-Ours 91.6 82.4
-IV-D Classification Performance on HAM10000
-To comprehensively evaluate the effectiveness of our neuromorphic skin lesion classification system, we conducted a series of controlled experiments in which a selection of state-of-the-art convolutional neural network (CNN) architectures were converted to spiking neural networks (SNNs) using our quantization-aware pipeline and executed on the same Akida hardware platform.
-This strategy ensures a fair comparison of all approaches under identical hardware constraints and SNN deployment settings.
-The evaluated models include both canonical CNN baselines and advanced architectures commonly used in medical image analysis.
-As shown in Table II, all evaluated CNN models show a decrease in accuracy and macro F1 score after conversion to SNN and deployment on neuromorphic hardware, with Top-1 accuracy values ranging from 83.2% to 88.1%.
-In contrast, our proposed model achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%, outperforming all baselines under the same SNN deployment conditions.
-This improvement demonstrates the benefit of our quantization-aware architecture and optimized network design for event-driven inference.
-The results confirm the practical utility of our method for real-time, resource-constrained medical applications, supporting efficient and accurate classification in portable diagnostic systems.
-TABLE III: Per-image inference latency and energy consumption of all models on the HAM10000 test set.
-CNN baselines are measured on an NVIDIA RTX 3090 (GPU) and Intel Xeon Gold 6226R (CPU); SNNs are measured on BrainChip Akida AKD1000.
+Ours 91.6 91.0
+4.2 Experimental Setup
+All model training, validation, and CNN-to-SNN conversion were conducted on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and NVIDIA RTX 3090 GPU, running Ubuntu 22.04 LTS.
+The software environment included Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1.
+Neuromorphic inference was performed on a BrainChip Akida AKD1000 PCIe board installed in the same system, with deployment and testing managed via the Akida Python API and default board settings.
+4.3 Analysis of Classification Results on HAM10000 Dataset
+Table 1 reports the class-wise precision, recall, and F1 score of QANA on the HAM10000 test set.
+Macro F1 is computed as the unweighted mean of per-class F1 scores on the test set.
+The model achieves consistent performance across lesion categories, including minority classes.
+With a Top-1 accuracy of 91.6% and a macro F1 of 91.0%, the system demonstrates effective discrimination of both common and rare lesions.
+Table 3: Per-image inference latency and energy consumption of all models on the HAM10000 test set.
+CNN baselines are measured on an NVIDIA RTX 3090 (GPU) and Intel i9-12900K (CPU); SNNs are measured on BrainChip Akida AKD1000.
 All values are averaged over 10,000 images.
-The rightmost columns show the percentage reduction achieved by SNNs on Akida compared to the corresponding CNN (GPU) version.
+The rightmost columns show the percentage reduction achieved by SNNs on Akida compared to the corresponding CNN version.
 Model CNN (GPU) CNN (CPU) SNN (Akida) Relative Reduction (%)
@@
 CNN Ensemble [1] 36.1 450.5 157.2 2177.1 7.5 8.6 79.2 98.1
-QANA (Ours) 27.6 163.1 83.9 841.5 1.5 1.7 94.6 98.6
-TABLE IV: Ablation study of core modules in our model on the HAM10000 test set. Metrics are reported as percentages (%). Each row shows the incremental addition of modules.
-Configuration Ghost Block ECA SE Quant. Head SMOTE Inc. Learn Accuracy Recall Precision F1 Score AUC-ROC
+QANA (Ours) 27.6 163.1 83.9 841.5 1.5 1.7 94.6 99.0
+Table 4: Ablation study of core modules in our model on the HAM10000 test set.
+Metrics are reported as percentages (%).
+Each row shows the incremental addition of modules.
+Inc.
+denotes fine-tuning of the readout layer after conversion using only validation data, while evaluation is performed on the fixed test split.
+Configuration Ghost ECA SE Quant. SMOTE Inc. Acc. Rec. Prec. F1 AUC
 Baseline 74.1 71.4 71.9 71.6 77.3
@@
 + SMOTE 91.0 89.2 90.0 89.6 92.7
-+ Incremental Learning 91.6 90.7 91.2 90.9 93.4
-IV-E Inference Speed and Energy Consumption
-We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures.
-All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA RTX 3090 GPU and an Intel Xeon CPU.
-For each model, the reported inference latency corresponds to the average per-image processing time over 1000 test samples.
-Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
-To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida.
-Table III presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations.
-Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures.
-When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 98.6% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
-For images requiring higher spatial detail, the $\times2$ tiling fallback increases latency and energy approximately in proportion to the number of patches while remaining within the power envelope of the edge platform.
-This fallback is applied conditionally based on a confidence threshold on the single-pass prediction and does not change the training protocol.
-TABLE V: Performance comparison of converted SNN models on the Clinical Dataset.
++ Incremental Learning 91.6 90.7 91.2 91.0 93.4
+4.4 Classification Performance on HAM10000
+To evaluate neuromorphic skin lesion classification under consistent deployment constraints, we trained a set of widely used CNN backbones and converted each model to an SNN using the same quantization-aware pipeline, then executed all converted networks on the same Akida platform with identical preprocessing and runtime settings.
+Table 2 summarizes the resulting Akida-SNN performance across canonical and advanced baselines, which typically exhibit reduced accuracy and macro F1 after conversion, with Top-1 accuracy spanning 83.2% to 88.1%.
+QANA attains 91.6% Top-1 accuracy and 91.0% macro F1 on HAM10000, yielding the best overall results under the same neuromorphic execution setting.
+Table 5: Performance comparison of converted SNN models on the Clinical Dataset.
 Model (SNN, Akida) Top-1 Accuracy (%) Macro F1 (%)
@@
 Ours 90.8 81.7
-IV-F Ablation Study of Model Components
-Table IV summarizes a stepwise ablation in which we start from a compact baseline and incrementally enable Ghost, ECA, SE, quantization-aware head with the same augmentation policy, embedding-space SMOTE, and on-chip incremental learning.
-All experiments use the same input resolution, optimizer, schedules, and 8-bit quantization configuration to isolate the effect of each component.
-Metrics are computed on the same test split.
-The drop observed for “Baseline + Ghost” is expected under a strictly width-preserving replacement of standard convolutions by Ghost modules.
-Ghost generates a fraction of channels through inexpensive linear transforms of a reduced set of primary channels.
-When used alone and early in the stack, this reduces the number of learned base filters that directly capture fine dermatoscopic structures such as pigment networks, streaks, and globules.
-The cheap transforms introduce correlated feature maps that do not add discriminative power without a mechanism to suppress redundancy.
-Under 8-bit activation clipping, these correlated responses compress the effective dynamic range and make the batch-norm folding more sensitive to local contrast, which in turn degrades minority-class recall.
-The combination of early downsampling and reduced primary channels further limits high-frequency detail preserved from $64\times 64$ inputs.
-As a result, the test accuracy decreases relative to the baseline when Ghost is introduced in isolation.
-Adding ECA recovers and surpasses the baseline by providing lightweight channel recalibration that suppresses redundant ghosted maps and reweights informative channels.
-The subsequent SE block contributes a complementary global channel prior that stabilizes inter-class boundaries and improves calibration.
-Enabling the quantization-aware head with the same augmentation policy improves the match between integer ranges at the readout and the upstream bounded activations, which tightens the loss landscape after BN folding.
-Injecting SMOTE in the embedding space raises coverage of minority manifolds without perturbing low-level image statistics and yields consistent gains in macro metrics.
-Finally, on-chip incremental learning adapts the last-stage parameters to post-conversion activation statistics and to the deployment data distribution, producing the highest accuracy and F1.
-We also observed that widening the primary branch of Ghost or delaying its use to later stages can mitigate the initial drop, but we kept the compact configuration to satisfy on-chip memory and latency constraints and relied on attention and calibration to recover accuracy.
-IV-G Classification Performance on Clinical Dataset
+4.5 Inference Speed and Energy Consumption
+We quantitatively evaluated inference latency and energy consumption against conventional CNN baselines and other converted SNN architectures.
+All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on the NVIDIA RTX 3090 GPU and the Intel Core i9-12900K CPU.
+Latency is reported for batch size $1$ with a fixed input pipeline, after a warm-up phase, and averaged over 10,000 images.
+Energy per image is computed as the mean device power during inference multiplied by the average per-image latency, using on-board power monitoring for Akida, NVIDIA-smi for GPU, and Intel RAPL for CPU.
+For each model, the reported inference latency corresponds to the average per-image processing time over 10,000 test samples.
+Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
+To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida.
+Table 3 presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations.
+QANA achieves the lowest inference latency and energy consumption among the evaluated architectures.
+When deployed as an SNN on Akida, it processes an image in 1.5 ms and consumes 1.7 mJ, which corresponds to a 94.6% latency reduction and a 99.0% energy reduction relative to its CNN implementation on the RTX 3090.
+4.6 Ablation Study of Model Components
+Table 4 reports the ablation results on the HAM10000 test set, illustrating the contribution of each core module to the overall performance.
+Modules were incrementally enabled to measure their isolated and cumulative effects on classification metrics, including accuracy, recall, precision, F1 score, and AUC-ROC (%).
+The results demonstrate that each module provides consistent improvements, with the complete model achieving the highest accuracy and F1 score.
+4.7 Classification Performance on Clinical Dataset
 We evaluated the proposed neuromorphic skin lesion classification system and several representative CNN architectures, all converted to SNNs and deployed on the Akida hardware platform.
-Table V presents the Top-1 accuracy and macro F1 score for each model under the same deployment conditions.
+Table 5 presents the Top-1 accuracy and macro F1 score for each model under the same deployment conditions.
 Our model achieves the highest accuracy and macro F1 score among all tested methods, confirming its effectiveness and robustness for neuromorphic inference in clinical scenarios.
-V Conclusion
-We presented QANA, a quantization-aware neuromorphic pipeline for skin lesion classification on edge hardware.
-Under a fixed $64\times 64$ input and identical preprocessing across methods, QANA reaches 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical dataset.
-The compiled SNN runs at 1.5 ms on chip and 1.7 mJ per image on Akida, measured as power above idle.
-These results indicate that conversion-preserving design and deployment-aware training enable accurate and efficient on-device dermatology inference.
-VI Acknowledgement
-This research was supported by the National Natural Science Foundation of China (Nos.
-62172336 and 62032018).
-The authors gratefully acknowledge BrainChip Holdings Ltd.
-for providing technical support and the Akida AKD1000 hardware platform, whose powerful neuromorphic computing capabilities enabled strong performance of the SNN model.
-The authors also extend their appreciation to Dr.
-Atif Mansoor and his teams for their preliminary contributions to this research.
+5 Conclusion
+This paper presented QANA, a quantization-aware neuromorphic framework for on-device skin lesion classification.
+Experiments on HAM10000 and a real-world clinical dataset show that QANA maintains competitive diagnostic accuracy while enabling efficient SNN deployment on BrainChip Akida.
+On Akida hardware, the converted model achieves 1.5 ms per-image latency and 1.7 mJ per image, supporting real-time inference under strict energy and privacy constraints.